### v2

# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

Abstract
A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communications (ISAC) system is proposed to enhance multi-target sensing and multi-user communication. Unlike the conventional RIS, the fRIS incorporates movable elements whose positions can be flexibly adjusted to provide extra spatial degrees of freedom. In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and communication symbol estimation error by optimizing the symbol estimator, transmit beamformer, fRIS phase shifts, and element positions. To solve this problem, an algorithm based on alternating minimization is devised, where subproblems are solved leveraging augmented Lagrangian method, quadratic programming, semidefinite-relaxation, and majorization-minimization techniques. A key challenge exists that the fRIS element positions affect both the incident and reflective channels, leading to the high-order composite functions regarding the positions. As a remedy, it is proved that the high-order terms can be transformed to linear and linear-difference forms using the characteristics of fRIS and structural channels, which facilitates the position optimization. Numerical results validate the effectiveness of the proposed scheme as compared to the conventional RIS-aided ISAC systems and other benchmarks.
Index Terms:
Alternating minimization, fluid reconfigurable intelligent surface, integrated sensing and communications, mismatch and estimation error
I Introduction
In the era of sixth-generation (6G) wireless communications, integrated sensing and communications (ISAC) emerged as a promising paradigm [1], addressing spectrum scarcity and reducing resource consumption. The ISAC aims to incorporate the sensing systems and the communication systems at different levels. Herein, the two systems shared the same spectrum and coexisted with minimal mutual interference [2] using null space projection [3], transceiver design [4], etc. To further mitigate interference, collaborations between the two systems were achieved by sharing critical information    [5], [6]. Additionally, considering the similarities in hardware and signal processing, the recent ISAC advancement focuses on designing unified platforms that can perform the dual functions [7]. For example, in [8], the transmit beamformer was designed by combining weighted radar waveforms and communication symbols. Besides, a wideband scenario about joint design was considered in [9], where mutual information was utilized as a unified metric to design the ISAC waveform.
Meanwhile, reconfigurable intelligent surface (RIS) garnered great attention in the 6G innovation. Specifically, the RIS consists of numerous reflective elements, each of which is able to induce phase shifts to the incident signals. This enables the RIS to dynamically adapt the channel propagation environment by designing the phase shifts [10]. Leveraging this characteristic, work [11] demonstrated significant improvement in communication sum-rate in RIS-aided multi-user scenarios. Apart from this, it was also shown that the RIS can suppress the undesired signals, such as eavesdropping  [12] and jamming [13]. Additionally, the RIS was applied in other diverse scenarios, including radar detection [14], near-field communications [15], and wireless power transfer and communications [16].
To harness the combined benefits of ISAC and RIS, extensive research was conducted on RIS-aided ISAC systems [17]. In [18], authors investigated an echo signal-to-noise ratio (SNR) maximization problem in a RIS-aided ISAC system, which serves a communication user and tracks a target through joint design of the transmit precoder and RIS beamformer. Building on this, [19] extended this framework to multi-user and multi-target scenarios under practical constraints. Furthermore, authors in [20] and [21] considered the energy efficiency maximization for the RIS-aided ISAC systems without compromising the communication and sensing performance. Additionally, researchers examined various other applications of the RIS-aided ISAC systems, such as simultaneously transmitting and reflecting surface [22], unsupervised learning [23], and extended reality [24].
While the RIS-aided ISAC systems offer significant enhancements for 6G communications, there remains room for the improvement in spatial degrees of freedom (DoFs) of the systems and the reduction in system complexity. Particularly, the antennas of the conventional arrays are fixed after assembly, which restricts the spatial diversity [25]. On the other hand, to achieve excellent performance, a massive number of elements are required, which increases manufacturing costs and exacerbates high-dimensional optimization complexity. To address these limitations, movable antenna, also known as fluid antenna, becomes a promising candidate. Enabled by the advancement of software-controllable fluidic metamaterials, the movable antenna can dynamically adjust its position in real time, offering additional spatial DoFs  [26], [27]. Moreover, the enhanced performance brought by the movable antenna allows comparable gains with fewer elements, significantly reducing the computational complexity caused by high-dimensional optimization.
The potential of the fluid antenna system was extensively studied. For instance, researchers demonstrated that it can promote the capacity improvement in point-to-point communication systems [25] and enhance downlink sum-rate in multi-user communication scenarios [28]. Besides, it was applied in ISAC systems to jointly optimize transceiver designs and antenna positions, enhancing communication and sensing performance [29], [30], [31]. Meanwhile, efforts to integrate the RIS with the fluid antenna system was made, where authors in [32] and [33] analyzed the outage probability bounds and asymptotic distributions for fluid antenna-aided RIS systems. Nevertheless, these works still rely on fixed RIS elements and do not fully exploit the spatial DoFs of RIS elements. By contrast, [34] introduced movable RIS elements with flexible positioning and studied the elimination of phase distribution offset in the system. However, this work was limited to a simple single-antenna and single-user communication setup, while did not consider the system optimization problem. This leaves room for further exploration in the potential of the movable RIS elements in ISAC systems.
Against the aforementioned background, we are motivated to study an ISAC system aided by a fluid RIS (fRIS). In particular, fRIS, named after the fluid antenna system, implies that the elements of RIS are moveable. In this system, beams are formed at the fRIS to serve the communication users and illuminate the targets, where the sensing beampattern mismatch and communication decoding error are jointly minimized. This problem is solved by an alternating minimization (AM)-based algorithm, which incorporates augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite-relaxation (SDR), and majorization-minimization (MM) technique. The main contributions of this work are summarized as:
  • As the movable elements can provide additional spatial DoFs compared to the conventional RIS, we propose a novel fRIS-aided ISAC system to improve the communication and sensing performance. Specifically, the ISAC transmit signal, fRIS beampattern for target illumination, communication symbol decoding and fRIS-related channels are characterized, and a joint sensing beampattern mismatch and communication decoding error minimization problem is formulated.
  • To tackle the resultant optimization problem, we devise an algorithm based on AM. First, we reformulate the original high-order problem into a more manageable form. Then, the problem is divided into several subproblems and dedicated optimization approaches are established for each subproblem. Eventually, the convergence and the computational complexity of the proposed algorithm are analyzed.
  • In the subproblem of optimizing the fRIS element positions, the optimization process is challenging since the changing RIS elements simultaneously affects the incident channel, the reflective channels. To address this, we leverage the characteristics of the fRIS and the structural channels to reduce the fourth-order terms of the element position to the linear and linear difference terms, so that the objective function of the subproblem can be simplified. This facilitates the optimization process.
  • Numerical results are demonstrated to validate the effectiveness of the proposed model. The results reveal a significant performance gain achieved by employing the fRIS in the ISAC system and the proposed scheme outperforms the benchmarks. Moreover, by employing the fRIS, the system performance loss caused by the reduction of the element number can be complemented, which alleviates the curse of dimensionality in the RIS optimization.
This paper is organized as follows. Section II introduces the model architecture of the fRIS-aided ISAC system, while formulates the symbol error and beampattern mismatch minimization problem. In section III, the AM-based algorithm for the resultant problem and the analysis of the proposed algorithm are presented in details. Section IV demonstrates the numerical results. Finally, conclusions are drawn in Section V.
[…]
II System Model and Problem Formulation
II-A System Model
Figure 1: A scenario of a fRIS-aided ISAC system that performs multi-target sensing and multi-user communication.
As depicted in Fig.1, we consider an ISAC system aided by a fRIS. Specifically, the ISAC base station (BS), equipped with a $M$-antenna uniform linear array (ULA), aims to simultaneously serve $K$ single-antenna communication users and sense $T$ point targets. Without loss of generalization, it is considered that the direct sensing and communication paths are blocked so that the fRIS is deployed to establish the virtual links and sensing paths. The fRIS has $N$ movable elements, where all the elements can move freely within a planar region $\mathcal{A}$ with the shape of $A\times A$. For $n$-th fRIS element, its position is denoted by its Cartesian coordinate $\mathbf{p}_{n}=\left[p_{x,n},p_{y,n}\right]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$. As a result, the positions of all the elements can be given compactly as $\mathbf{p}=\left[\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}\right]\in\mathcal{A}^{2\times N}$.
In this system, we denote the data symbols for the users as $\mathbf{s}_{c}\in\mathbb{C}^{K\times 1}$, which is precoded by a beamformer $\mathbf{W}\in\mathbb{C}^{M\times K}$ in the BS. This yields the precoded digital signal as
$\mathbf{x}=\mathbf{W}\mathbf{s}_{c}\in\mathbb{C}^{M\times 1}.$ (1)
The symbol $\mathbf{s}_{c}$ is considered statistically independent and auto-correlated, namely $\mathbb{E}[\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}]=\mathbf{I}_{K\times K}$. After converting the radio frequency domain, the signal $\mathbf{x}$ is emitted from the BS and propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS. The phase of the incident signal in the fRIS is altered adaptively by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\left\{\theta_{n},\forall n=1,\dots,N\right\}$ represents the phase shift coefficient for $n$-th RIS element. As a result, the reflective signal at the fRIS can be given as
$\mathbf{v}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}.$ (2)
Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and propagate through the reflecting channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ to serve the communication users.
II-A1 Sensing Model
As we know, the illumination power to the targets is direction related to the echo SNR, where more target information can obtained if the echo SNR is high. As a consequence, beams should be formed at the fRIS towards the directions of the targets to illuminate the targets. To this end, one can design the correlation matrix of the reflective signal $\mathbf{v}$ to manipulate the beampattern of the reflective signal. Specifically, the relationship between the beampattern and the correlation matrix of the reflective signal can be expressed as
$\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ $=\mathbf{a}^{\mathrm{H}}(\phi,\psi,\mathbf{p})\mathbf{R}_{s}\mathbf{a}(\phi,\psi,\mathbf{p}),$ (3)
in which $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$ is the covariance matrix of signal $\mathbf{v}$. The symbol $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS, which is not only dependent on azimuth angle $\phi$ and elevate angle $\psi$, but also the positions of the movable elements $\mathbf{p}$. For conciseness, we leave the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in the subsection II-A3.
For beam design, a common metric is the beampattern similarity, where the mismatch between the practical beampattern $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and the ideal pattern $\mathcal{P}_{d}(\phi,\psi)$ are measured to evaluate the target sensing performance. Specifically, the beampattern mismatch can be written as
$\varepsilon_{r}=\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2},$ (4)
in which $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote the grid number of the azimuth angles and elevate angles, respectively. When $\varepsilon_{r}$ is minimized, the practical beampattern aligns with the ideal beampattern, such that the reflective signal can effectively illuminate the targets.
II-A2 Communication Model
Apart from illuminating the targets, the reflective signal $\mathbf{v}$ also propagates to the users for communication through the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$. The signal received at user $k$ can be given as
$y_{c,k}=\mathbf{h}_{rc,k}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+n_{c,k},~{}~{}\forall k=1,\cdots,K,$ (5)
where $n_{c,k}$ represents the additive Gaussian white noise at the receiver of user $k$ with zero mean and $\sigma_{0}^{2}$ variance. To estimate the transmitted symbols from the received signal, an estimation factor $\omega$ is introduced, i.e.,
$\hat{s}_{c,k}=\omega y_{c,k}.$ (6)
Given the estimated communication symbols, a direct performance indicator for evaluating the communication performance is the mean square error (MSE) between the estimated symbol and the transmitted symbol. It can be formulated as
$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$ (7a)
=$\overset{(a)}{=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (7b)
where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ incorporates all the reflective channels. The operation $(a)$ substitutes (5) and (6) to (7a) and leverages the statistical independence and the auto-correlation characteristics of $\mathbf{s}_{c}$.
II-A3 Channel Model
As stated before, the steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$ as well as the azimuth angle $\phi$ and the elevate angle $\psi$. Specifically, the steering vector of the fRIS can be expressed as [25]:
$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!.$ (8)
The symbol $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, which gives
$d_{n}(\phi,\psi,\mathbf{p}_{n})=p_{x,n}\sin(\phi)\cos(\psi)+p_{y,n}\sin(\psi).$ (9)
Without loss of generalization, the fRIS is typically deployed in the area where line-of-sight (LoS) paths dominate, so that the signal coverage can be extended. Consequently, the fRIS-related channels are modeled as LoS channels, which are determined by the steering vector of the fRIS. For the channel $\mathbf{G}$ from the BS to the fRIS, it can be modeled as
$\mathbf{G}$ $=\sqrt{\zeta_{G}}\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}(\phi_{t}),$ (10)
in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth direction and $\psi_{r}$ in elevate direction, which is in the form of (8). On the other hand, the ULA is equipped in the BS, so that the transmit steering vector $\mathbf{a}_{t}$ in the direction of $\phi_{t}$ can be written as
$\mathbf{a}_{t}=\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{1}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{m}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{M-1}}\end{bmatrix}^{\mathrm{T}},$ (11)
where $\{\widetilde{d}_{m}=\frac{\lambda}{2}m\sin(\phi_{t}),~{}\forall m=0,\cdots,M-1\}$ and the inter-element spacing is half wavelength. Similarly, the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}$ from the fRIS to the user $k$ is modeled as
$\mathbf{h}_{rc,k}^{\mathrm{H}}=\sqrt{\zeta_{k}}\mathbf{a}^{\mathrm{H}}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p}),\forall k=1,\cdots,K,$ (12)
where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the transmit steering vector of the fRIS to user $k$ from azimuth direction $\phi_{c,k}$ and elevate direction $\psi_{c,k}$, which takes the form of (8). The path loss between the fRIS and the user $k$ is denoted as $\sqrt{\zeta_{k}}$. Accordingly, the overall reflective channel $\mathbf{H}_{rc}$ can be rewritten as
$\mathbf{H}_{rc}$ $=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}=\mathbf{A}_{rc}\boldsymbol{\Sigma}_{rc},$ (13)
where we denote
$\mathbf{A}_{rc}$ $=\begin{bmatrix}\mathbf{a}_{c,1}(\phi_{c,1},\psi_{c,1},\mathbf{p})&\!\!\!\cdots\!\!\!&\mathbf{a}_{c,K}(\phi_{c,K},\psi_{c,K},\mathbf{p})\end{bmatrix},$ (14a)
$\boldsymbol{\Sigma}_{rc}$ $=\mathrm{Diag}\left(\begin{bmatrix}\sqrt{\zeta_{1}}&\cdots&\sqrt{\zeta_{K}}\end{bmatrix}\right).$ (14b)
II-B Problem Formulation
In order to simultaneously achieve the sensing and the communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (4) and the communication symbol estimation MSE $\varepsilon_{c}$ of (7) is minimized by optimizing the transmit beamformer $\mathbf{W}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ as well as the scaling factor $\beta$ of the sensing beampattern. Consequently, the problem can be formulated as
$\min_{\mathbf{W},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~{}\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
$+(1\!-\!\alpha)\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\right)$ (15a)
$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (15b)
$~{}\omega\in\mathbb{R},$ (15c)
$~{}|\boldsymbol{\Theta}|_{n,n}=1,\forall n=1,\cdots,N$ (15d)
$~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N$ (15e)
$~{}\|\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}\!\geq\!\Delta D,\forall n,n^{\prime}\!\!=\!\!1,\!\cdots\!,N,n\!\neq\!n^{\prime},$ (15f)
where $\alpha$ is the weighting coefficient indicating the bias of the sensing and the communication. The constraint (15b) represents that the transmit power should not exceed $P_{t}$, while in the constraint (15d), the modulus-one phase shift of the fRIS is required. The constraints (15e) and (15f) indicate that each RIS element ought to move within the pre-set planar region, and the inter-space among elements should be larger than $\Delta D$. It is challenging to solve the problem (15) owing to the two reasons. On the one hand, (15d) is a constant modulus constraint, while (15f) takes the non-convex form of “Norm $\geq$ Constant”. Besides, the objective function is highly complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to different variables.
III Proposed Algorithm
In this section, an effective AM-based algorithm is developed to address the problem (15) by iteratively updating the variables, and we analyze the convergence and computational complexity of the proposed algorithm.
III-A Reformulation of $\varepsilon_{r}$
Carefully inspecting $\varepsilon_{r}$ in (4), it can be observed that $\varepsilon_{r}$ is a quadratic function with respect to the covariance matrix $\mathbf{R}_{s}$, while $\mathbf{R}_{s}$ is also a quadratic function with respect to the variables, making $\varepsilon_{r}$ a quartic function with respect to the variables. To simplify the objective function in (15a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function with respect to the variables.
Specifically, we first design an optimum covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only takes the sensing into consideration. According to [35], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can given as
$\min_{\beta,\widetilde{\mathbf{R}}_{s}}$ $~{}\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}},\widetilde{\mathbf{R}}_{s})\right|^{2}$ (16a)
$\mathrm{s.t.}$ $~{}\beta>0,\widetilde{\mathbf{R}}_{s}\succeq 0,$ (16b)
$\mathrm{tr}\left(\widetilde{\mathbf{R}}_{s}\right)={\|\mathbf{G}\mathbf{x}\|^{2}},$ (16c)
$\mathrm{rank}\left(\widetilde{\mathbf{R}}_{s}\right)=1,$ (16d)
where the objective function requires the designed sensing beampattern to be similar to an ideal one, the constraint (16b) indicates that $\widetilde{\mathbf{R}}_{s}$ is semi-definite, the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$ in the constraint (16c), and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since the ultimately designed $\mathbf{R}_{s}$ should satisfy $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$. The problem (16) can be directly solved with SDR approach. To guarantee the rank-one constraint of $\widetilde{\mathbf{R}}_{s}$, various methods can be employed, such as Gaussian randomization [36] and penalty-based method [37].
With the obtained $\widetilde{\mathbf{R}}_{s}$ designed in (16), the beams can be effectively formed towards the targets. Considering that the designed $\widetilde{\mathbf{R}}_{s}$ is rank-one, the corresponding reflective signal at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be retrieved by the performing eigenvalue decomposition to $\widetilde{\mathbf{R}}_{s}$. This yields
$\widetilde{\mathbf{R}}_{s}=\mathbf{s}_{r}\mathbf{s}_{r}^{\mathrm{H}},$ (17)
where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$, while $\lambda_{s}$ and $\mathbf{u}_{s}$ represent the maximum eigenvalue and the corresponding eigenvector of $\widetilde{\mathbf{R}}_{s}$, respectively. The factorized signal $\mathbf{s}_{r}$ represents an optimum and achievable reflective signal whose correlation matrix is $\widetilde{\mathbf{R}}_{s}$ producing directive beams towards the targets. In other words, $\mathbf{s}_{r}$ can be viewed as a desired reflective signal for sensing the targets.
Consequently, the sensing metric is reformulated as MSE between the practical reflective signal $\mathbf{v}$ at the fRIS and the desired reflective signal $\mathbf{s}_{r}$, i.e.
$\widetilde{\varepsilon}_{r}=\|\mathbf{s}_{r}-\mathbf{v}\|^{2}=\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (18)
It can be observed that the reformulated sensing metric $\widetilde{\varepsilon}_{r}$ becomes quadratic with respect to all the variables, which simplifies the subsequent optimization.
Ultimately, the problem (15) can be reformulated as
$\min_{\mathbf{W},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~{}\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
$+(1\!-\!\alpha)\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\right)$ (19a)
[…]
However, the problem is still challenging owing to the non-convex constraints and objective function. In the subsequence, the problem (19) is solved by iteratively updating each variable with others being fixed. It is worth mentioning that $\mathbf{s}_{r}$ may change after the update of $\mathbf{p}$, but according to the Lyapunov stability [38], $\mathbf{s}_{r}$ is asymptotically stable over iterations due to the bounded $\mathbf{p}$ and the finite resources.
III-B Optimization of Communication Symbol Estimator $\omega$
When $\mathbf{x}$, $\mathbf{p}$ and $\boldsymbol{\Theta}$ are fixed, the objective function in the problem (19) is a quadratic function with respect to $\omega$. After ignoring the terms that are unrelated to $\omega$, the problem can be simplified as
$\min_{\omega\in\mathbb{R}}$ $~{}~{}\varepsilon_{c}=\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2}.$ (20)
The optimal $\omega^{\star}$ can therefore be derived by letting the derivative of $\varepsilon_{c}$ be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$. This yields a closed-form optimum:
$\omega^{\star}=\frac{\mathfrak{R}\{\mathbf{s}^{\mathrm{H}}_{c}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}}{\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+K\sigma_{0}^{2}}.$ (21)
III-C Optimization of fRIS Phase Shift $\boldsymbol{\Theta}$
Given $\mathbf{W}$, $\mathbf{p}$ and $\omega$, the problem (19) can be reduced to a constant-modulus constrained quadratic minimization problem, which is given as
$\min_{\boldsymbol{\Theta}}$ $~{}\alpha\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!(1\!-\!\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (22a)
$\mathrm{s.t.}$ $~{}|\boldsymbol{\Theta}|_{n,n}=1,~{}~{}\forall~{}n=1,\cdots,N.$ (22b)
The challenge of this problem mainly stems from the non-convex constant-modulus constraint. To address this task, one can first let $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [21] to transform the problem (22) to a compact form:
$\min_{\boldsymbol{\theta}}$ $~{}\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{A}}_{1}\boldsymbol{\theta}-\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{b}}_{1}-{\mathbf{b}}_{1}^{\mathrm{H}}\boldsymbol{\theta}$ (23a)
$\mathrm{s.t.}$ $~{}|\boldsymbol{\theta}|_{n}=1,\forall n=1,\cdots,N.$ (23b)
where
${\mathbf{A}}_{1}$ $\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]$
$~{}~{}+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right],$ (24a)
${\mathbf{b}}_{1}$ $=\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})+\omega(1-\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}).$ (24b)
The nonconvex constant-modulus constraint (23b) remains challenging, where various approaches have been developed in the wide literature to effectively handle such a constraint, such as SDR [39] and manifold optimization [40]. Therefore, the details will not be elaborated upon here.
III-D Optimization of Transmit Beamformer $\mathbf{W}$
In this subsection, the optimization of the transmit beamformer $\mathbf{W}$ will be discussed. Viewing other variables as constant, the problem (19) can be reduced to
$\min_{\mathbf{W}}$ $~{}\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)$ (25a)
$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (25b)
where we denote $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ as the overall channel from the transmitter to the users. For the objective function (25a), the norm operations are expanded and thus it becomes
$\alpha\!\left(\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\}\right)$
$+$ $(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\}\right).$ (26)
By leveraging the cyclic property of trace, the relationship of $\mathrm{tr}(\mathbf{X}^{\mathrm{H}}\mathbf{C}_{1}\mathbf{X}\mathbf{C}_{2})=\mathrm{vec}^{\mathrm{H}}(\mathbf{X})(\mathbf{C}_{2}^{\mathrm{T}}\otimes\mathbf{C}_{1})\mathrm{vec}(\mathbf{X})$ and $\mathrm{tr}(\mathbf{C}^{\mathrm{H}}\mathbf{X})=\mathrm{vec}^{\mathrm{H}}(\mathbf{C})\mathrm{vec}(\mathbf{X})$, the objective function can further be simplified as
$\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\},$ (27)
in which
$\mathbf{w}$ $=\mathrm{vec}(\mathbf{W}),$ (28a)
$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{G}^{\mathrm{H}}\mathbf{G}\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right),$ (28b)
$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\mathbf{s}_{c}^{\mathrm{H}}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}\right).$ (28c)
It is worth noting that in (28b), $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ has been simplified as $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ since $\boldsymbol{\Theta}$ is a modulus-one diagonal matrix and $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ always holds. After this transformation, it can be seen that the objective (27) is a convex quadratic function. As for the constraint (25b), it can also be re-expressed as the inner product of $\mathbf{w}$, that is, $\mathbf{w}^{\mathrm{H}}\mathbf{w}\leq P_{t}$. In this context, we consider full power usage in the transmitter, namely $\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t}$. As a result, the problem (25) can be written as
$\min_{\mathbf{w}}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\}$ (29a)
$\mathrm{s.t.}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t},$ (29b)
which can be addressed by ALM [41, 42]. Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part. The problem can thus be converted to its real form as
$\min_{\overline{\mathbf{w}}}$ $~{}\mathring{f}(\overline{\mathbf{w}})=\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{w}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{w}}$ (30a)
$\mathrm{s.t.}$ $~{}\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}=P_{t},$ (30b)
where
$\overline{\mathbf{w}}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{w}\}\\ \mathfrak{I}\{\mathbf{w}\}\end{bmatrix},\overline{\mathbf{b}}_{2}=\begin{bmatrix}\mathfrak{R}\{\mathbf{b}_{2}\}\\ \mathfrak{I}\{\mathbf{b}_{2}\}\end{bmatrix},$ (31a)
$\overline{\mathbf{A}}_{2}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{A}_{2}\}&-\mathfrak{I}\{\mathbf{A}_{2}\}\\ \mathfrak{I}\{\mathbf{A}_{2}\}&\mathfrak{R}\{\mathbf{A}_{2}\}\end{bmatrix}.$ (31b)
Accordingly, the augmented Lagrangian function of problem (30) can be written as
$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{w}})+\mu\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)^{2},$ (32)
in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. This can be tackled by the following iteration procedure:
$\overline{\mathbf{w}}^{(t)}$ $=\arg\min_{\overline{\mathbf{w}}}\mathcal{L}(\overline{\mathbf{w}},\mu^{(t-1)}),$ (33a)
$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{w}}^{\mathrm{T}(t)}\overline{\mathbf{w}}^{(t)}-P_{t}\right).$ (33b)
Specifically, one can solve the subproblem (33a) by quasi-Newton method [42], which is implemented by the fminunc tool. This process is terminated when $\left|\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right|<\epsilon_{0}$ is satisfied. It is noted that the penalty factor $\gamma$ gradually increases among the iterations for fast convergence. After obtaining $\overline{\mathbf{w}}^{\star}$, it is converted to the complex matrix form and reshaped as $\mathbf{W}^{\star}$.
III-E Optimization of fRIS Element Positions $\mathbf{p}$
In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$ with other variables being fixed. First, the problem (19) can be simplified by omitting the unrelated terms and constraints:
$\min_{\mathbf{p}}$ $~{}\alpha\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$
$~{}+(1-\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$ (34a)
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N,$ (34b)
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}\geq\Delta D,\forall n,n^{\prime}=1,\cdots,N,n\neq n^{\prime}.$ (34c)
As mentioned in section II-A3, the channel $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the fRIS element positions. It is challenging to solve the problem (34) owing to two main reasons. First, the objective function is a complicated composite function concerning the variable $\mathbf{p}$. Since the fRIS element positions affect both the incident channel $\mathbf{G}$ and the reflective channel $\mathbf{H}_{rc}$, there appear quadratic and quartic forms simultaneously in the outer function. In the interior function, the variable $\mathbf{p}$ is contained in the complex exponential function of an array manifold, shown in (8). On the other hand, the constraint (34c) takes the form of “Norm $\geq$ Constant”, rendering it non-convex.
To explicitly observe $\mathbf{p}$ in (34a), we process the outer function by substituting $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) to (34a) and integrating the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
Proposition 1. After omitting the constants, the objective function (34a) can be rewritten as
$\widetilde{f}_{0}=$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\right\}$
$+(1\!-\!\alpha)\left[c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\right.$
$\left.-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\}\right],$ (35)
where we denote $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$, $c_{0}=\omega^{2}\zeta_{G}c_{x}$, $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$, $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$, $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
Proof: Detailed derivations are provided in Appendix A.
However, the variable in (III-E) is contained in the array manifold, while there still exist $K$ quartic function terms, a quadratic function term and a linear function item with respect to the array manifold. Fortunately, it can be proved in the following proposition that the quartic, quadratic and linear items in (III-E) can be reduced to the first-order items and the difference of the first-order items if we only focus on optimizing the position of one element at a time.
Proposition 2. When considering the position of $n$-th element $\mathbf{p}_{n}$ while fixing others, the objective function $\widetilde{f}_{0}$ in (III-E) can be further simplified as
$\widetilde{f}_{1}(\mathbf{p}_{n})$ $=\widetilde{f}_{1,1}+\widetilde{f}_{1,2}+\widetilde{f}_{1,3}$
$=-\nu_{n}\cos\left(\xi_{n}+\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})\right)$
$~{}~{}+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
$~{}~{}-\sum_{k=1}^{K}\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (36)
In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of $n$-th entry of ${}\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant. The symbol $d_{r,n}(\mathbf{p}_{n})$ is $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$. In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$. Meanwhile, we denote $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$. In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$ and $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$.
Proof: Detailed derivations and symbol explanations are shown in Appendix B.
Observing the objective function $\widetilde{f}_{1}$, it contains cosine functions, which are neither convex nor concave with respect to $\mathbf{p}_{n}$. To address this problem, we employ the MM framework, where surrogate functions that locally upper bound the objective function $\widetilde{f}_{1}$ are constructed. Specifically, the constructions of the surrogate function is given by the following lemma.
Lemma 1. [43] Considering a function $\widetilde{f}$ with $\mathbf{x}$ being the variable, its second-order Taylor expansion at the given point $\mathbf{x}_{0}$ can be given as
$\widetilde{f}(\mathbf{x})\approx\widetilde{f}(\mathbf{x}_{0})$ $+(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})$
$+\frac{1}{2}(\mathbf{x}-\mathbf{x}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{x}_{0})(\mathbf{x}-\mathbf{x}_{0}),$ (37)
where $\nabla\widetilde{f}(\mathbf{x})$ is the gradient vector of $\widetilde{f}(\mathbf{x})$, while $\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$ represents the Hessian matrix of the function.
By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$, the function $\widetilde{f}(\mathbf{x})$ can be locally upper-bounded by
$\!\widetilde{f}(\mathbf{x})\leq\widetilde{f}(\mathbf{x}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{x}-\mathbf{x}_{0})\right\|^{2}.$ (38)
In particular, the parameter $\delta$ can be chosen as the Frobenius norm of the Hessian matrix $\nabla^{2}f(\mathbf{x}_{0})$.
According to the Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function. To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined. Specifically, the gradient and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ are given by the following theorem.
Theorem 1. According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which can be expressed as
$\nabla\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}}\end{bmatrix}^{\mathrm{T}}.$ (39)
For the Hessian matrix of $\widetilde{f}_{1}$, it can also be given as
$\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}\partial p_{y,n}}\\ \sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{y,n}}\end{bmatrix}.$ (40)
Specifically, there are 6 first-order derivative terms in (39) and 12 second-order derivative terms in (40), whose detailed expressions are left in the Appendix C for brevity.
Proof: Please refer to Appendix C.
According to Lemma 1 and Theorem 1, the equivalent objective function ultimately be given as
$(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\!+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2},$ (41)
where the constants are neglected and $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$. The $\mathbf{p}_{n}^{(t-1)}$ represents the output position of $n$-th element from the last round optimization. Apparently, the function (41) is a convex quadratic function with respect to $\mathbf{p}_{n}$.
Apart from the objective function, the constraint (34c) remains non-convex in the current form. To deal with it, an effective approach is to relax the constraint (34c) with the first-order Taylor expansion. Specifically, the first-order Taylor expansion of the left term in (34c) can be given as
$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}$
$~{}~{}~{}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42a)
$=\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42b)
As a result, the constraint (34c) can be relaxed as
$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n,n^{\prime}\!=\!1:N,n\!\neq\!n^{\prime},$ (43)
which is affine regarding $\mathbf{p}_{n}$. Therefore, the problem for optimizing the position of the $n$-th element is given as
$\min_{\mathbf{p}_{n}}$ $~{}(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2}$ (44a)
$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},$ (44b)
$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n\!\neq\!n^{\prime}.$ (44c)
The problem (44) is a typical convex quadratic problem, which can be solved by using quadprog [44, 45]. Besides, we adopt circle packing scheme [46] for the initialization of fRIS element positions, where the initial positions are sufficiently separated. The details of the proposed algorithm are summarized in Algorithm 1.
Algorithm 1 Optimization of fRIS element positions
0:    Channel parameters, transmit signal, fRIS phase shift matrix, system parameters, desired reflective signal and communication symbols;
1:  Initialize the element positions $\mathbf{p}^{(0)}$ with the circle packing scheme;
2:  for $n=1\longrightarrow N$ do
3:     Calculate the gradient and the Hessian matrix of $\widetilde{f}_{1}$ according to the Theorem 1;
4:     Calculate $\delta_{n}$ based on the Hessian matrix of $\widetilde{f}_{1}$;
5:     Solve the problem (44) using quadprog;
6:  end for
7:  return  fRIS element positions $\mathbf{p}$.
III-F Convergence and Computational Complexity Analysis
The proposed algorithm is summarized in Algorithm 2, which incorporates the solutions derived above. The algorithm iteratively optimizes the variables, where the process continues until the objective function value in (19) becomes stable.
III-F1 Convergence Analysis
In this part, we analyze that the convergence of the proposed Algorithm 2. First, the subproblem (20) is a unconstrained convex quadratic problem with a close-form solution, and thus $\omega^{t}$ update does not increase objective function value:
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (45)
Updating $\mathbf{\Theta}^{t}$ in (23), no matter utilizing the SDR [39] or the manifold method [40], it was proved that the objective function value of (23) is non-increasing over iterations, such that
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (46)
In the optimization of $\mathbf{W}$, quasi-Newton method guarantees that the update of $\overline{\mathbf{w}}$ is non-increasing [42], while equality constraint is satisfied if the penalty parameter becomes sufficiently large. As a result, the solution to (33) can satisfy the equality constraint, while the process (33) is non-increasing:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t}\!,\mathbf{p}^{t}).$ (47)
Proved in [25], the subproblem (44) is a non-decreasing process given the nature of its convexity and the inequalities of Taylor expansions, while the constrained optimization process can converge in the MM framework when the function is smooth objective function [47]:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t+1})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t}).$ (48)
Besides, the finite resources indicate that the objective function value is bounded and greater than zero. As a result, the Algorithm 2 can guarantee to converge.
III-F2 Computational Complexity Analysis
In this part, the computational complexity is analyzed.
  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving the problem (16) with SDR and the eigenvalue decomposition. Given a solution accuracy $\epsilon_{1}$, the computational complexity of SDR is $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$ [36], while the eigenvalue decomposition has the computational complexity of $\mathcal{O}\left(N^{3}\right)$.
  • The optimal symbol estimator can be calculated via a close-form expression, whose computational complexity stems from matrix multiplications. Obtaining the numerator in (21) has the computation complexity of $\mathcal{O}\left(KN+N^{2}+MN+M\right)$, while calculating the denominator has the computation complexity of $\mathcal{O}\left(2(MK+NK+N^{2})+M\right)$.
  • When optimize the fRIS phase shift, the computational complexity of calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$. Besides, the SDR approach has the computational complexity of $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{2})\right)$ with $\epsilon_{2}$ being the solution accuracy, while that of the manifold optimization is $\mathcal{O}\left(N^{1.5}\right)$ [40].
  • In the optimization of transmit beamformer design, the computational complexity of matrix multiplication is $\mathcal{O}\left(M^{2}K^{2}\!+\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!+\!K^{2}\right)$, while the design of the transmit beamformer with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$ with $\epsilon_{3}$ being the convergence accuracy [48, 49].
  • When optimize the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$. Additionally, the computational complexity of solving the problem (44) is $\mathcal{O}\left(\left(N^{1.5}\log(1/\epsilon_{4})\right)\right)$ [25] with accuracy $\epsilon_{4}$ for the interior-point method.
As a result, the total computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$, neglecting the lower-order terms.
Algorithm 2 Proposed alternating optimization algorithm
0:    Channel parameters, noise power, transmit power budget, ideal sensing beampattern, weighting coefficient, users number, fRIS planar region size, and the minimum inter-element spacing requirement.
1:  Initialize all the variables;
2:  while no convergence of $\varepsilon$ do
3:     Calculate correlation matrix for sensing-only by (16) and obtain desired reflective signal $\mathbf{s}_{r}$ via (17);
4:     Update communication symbol estimator by (21);
5:     Optimize $\boldsymbol{\Theta}$ by solving problem (23);
6:     Update ${\mathbf{W}}$ by implementing the ALM process in (33);
7:     Optimize fRIS element positions via Algorithm 1;
8:  end while
9:  return  Optimized variables $\mathbf{W}$, $\boldsymbol{\Theta}$, $\mathbf{p}$, $\omega$, $\beta$ and the objective function value.
IV Simulation Results
This section provides numerical results to validate the effectiveness of the fRIS-aided ISAC system and the proposed algorithm. Unless stated otherwise, the simulation parameters are set as follows. The center frequency of this system is $2.4$GHz, corresponding to a wavelength of $\lambda\!=\!0.125$m. The ISAC BS is located at ($3$m, 00m, 00m), and equipped with $M\!=\!8$ antennas. The transmitter has a maximum transmit power of $10$dBm. The fRIS, consisting of $16$ elements ($N_{x}\!=\!N_{y}=4$), is positioned at (00m, $3$m, $3$m), with a size of $A\!\times\!A$ where $A\!=\!4\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$. Additionally, we consider $K\!\!=\!\!4$ communication users and $T\!\!=\!\!3$ targets. The users are randomly distributed within an area centered at ($30$m, $100$m, 00m) with a radius of $10$m. As observed from fRIS, the targets’ (azimuth, elevate) angles are ($-60^{\circ}$, $0^{\circ}$), ($10^{\circ}$, $0^{\circ}$) and ($55^{\circ}$, $0^{\circ}$), respectively. The noise power of the users are set as $\sigma_{0}^{2}\!=\!-60$dBm and QPSK is used for modulation. The weighting coefficient $\alpha$ for balancing the communication and sensing performance is $0.5$. Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-10$dB is the path loss exponent and dist denotes the transmitter-receiver distance [39]. The penalty factor for ALM and the convergence threshold are set as $1$ and $10^{-5}$, respectively. All simulations are averaged over $100$ Monte Carlo trials.
Figure 2: Convergence behaviors of the proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step.
Figure 3: The effect of RIS element number: the objective function value $\varepsilon$ versus the element number.
The performance of the proposed design is evaluated against the following benchmark schemes:
  • Conventional RIS-aided ISAC system (Conven): The positions of RIS elements are fixed and uniformly distributed across the fRIS plane. The rest of the beamforming designs follow the same procedures as proposed.
  • Discrete position selection (DPS): The moving region of the elements is quantized into discrete grids of equal spacing, where greedy algorithm is employed for the position selection. Specifically, the position of each element is selected sequentially in the discrete position set, while the positions of the other elements are kept unchanged.
  • Random RIS element deployment (Rand): The positions of RIS elements remain fixed for each channel realization but randomly vary across realizations. The optimization of other components is performed as the proposed scheme.
Figure 4: Beampattern at the fRIS: beam gain (dB) versus spatial angle (degree). (a) The beampattern when bias to the communication ($\alpha=0.1$). (b) The beampattern when no bias ($\alpha=0.5$). (c) The beampattern when bias to the sensing ($\alpha=0.9$). (d) The beampattern of different approaches when $\alpha=0.5$.
TABLE I: Comparisons in terms of Average Performance Value
and Time Complexity per round
Scheme Objective Function Value Elapsed Time (s)
Proposed 0.378 313.52
($N=25$)
Conven 0.394 934.01
($N=64$)
Conven 0.432 372.46
($N=49$)
In Fig.2, we examine the convergence performance of our proposed algorithm for the fRIS-assisted ISAC system and compare it with the benchmark scenarios. As expected, the proposed approach demonstrates convergence, with the objective function value decreasing rapidly and stabilizing at $0.57$ after $15$ iterations. Meanwhile, it can be observed that the joint MSE of the proposed fRIS-aided ISAC case outperforms the baselines, especially compared to the conventional RIS-aided ISAC case. This improvement can be attributed to the continuous and flexible element position optimization, which provides additional DoFs in the spatial domain.
Fig.3 illustrates the relationship between the number of fRIS elements ($N$) and the objective function value ($\varepsilon$), with $N$ varying from $16$ to $64$. As observed, the objective function value of the proposed scheme and the baselines all decrease as $N$ increases, while the proposed scheme achieves the best performance among all. In addition, in Table I, we compare the average objective function value $\varepsilon$ and the time complexity of the proposed scheme with those of the conventional RIS-aided ISAC system. Specifically, we select the proposed scheme with $N=25$, the “Conven” scheme with $N=49$ and $N=64$, where either the objective function value or the elapsed time has similar performance. When compared with the “Conven” scheme with $N=64$, the proposed scheme with $N=25$ can achieve slightly better objective function value yet requiring $1/3$ of the elapsed time. When compared with the “Conven” scheme with $N=49$, the proposed scheme with $N=25$ exhibits comparable elapsed time but significantly better objective function value performance. These results demonstrate that employing the fRIS not only compensates for performance losses caused by reducing RIS elements but also alleviates the complexity of high-dimensional optimization.
To evaluate the system’s sensing performance, we further investigate the fRIS beampattern under different weighting coefficients, and compare with the ideal case and the baselines in Fig.4. It is worth noting that the targets are assumed to lie at an angle of $0^{\circ}$ in elevate, so that the beampatterns of the cross section at $0^{\circ}$ in elevate are plotted for clearer observations and easier comparisons. In Fig.4 - Fig.4, the practical, sensing-only and ideal beampatterns are illustrated for the weighted coefficients $\alpha=0.1,0.5\text{ and }0.9$, respectively. It can be observed that beams are consistently formed in the target directions across all $\alpha$ values. However, in the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels. In contrast, when $\alpha=0.5$ and $\alpha=0.9$ where the system is increasingly biased towards the sensing, the practical beams gradually align with the sensing-only case. On the other hand, Fig.4 compares the beampattern obtained from different approaches when $\alpha=0.5$. It can be found that all the beampatterns have formed peaks in the target directions, but the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the ”DPS” and ”Rand” cases.
Figure 5: Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients.
Figure 6: Average BER (dB) versus different receive noise power (dBm) with different approaches.
In addition, we evaluate the communication performance of the system under different settings. Fig.5 depicts the average bit error rate (BER) of the proposed scheme versus various noise power levels, ranging from $-40$dBm to $-80$dBm. As expected, the average BER decreases when the noise power at the receivers decreases, and QPSK modulation exhibits lower BER when compared to 16QAM. Moreover, when $\alpha$ decreases, the BER further improves as the system becomes more communication-biased. Notably, at a noise power of $-80$dBm and a weighting coefficient of $\alpha=0.9$, the BER can drop to approximately $-28$dB with QPSK modulation. On the other hand, Fig. 6 compares the average BER of the proposed method to the benchmarks. As shown, all cases exhibit a decreasing BER trend as noise power level reduces, while the proposed method consistently achieves a lower BER compared to the baselines.
Figure 7: The effect of the fRIS region size: the objective value versus the normalized region size $A/\lambda$.
In Fig.7, the impact of the normalized region size $A/\lambda$ on the proposed scheme is illustrated, where $A$ varies from $2\lambda$ to $6\lambda$. The results show that the objective function value $\varepsilon$ decreases as the region size increases, which is expected since the larger size of the fRIS provides more DoFs for the mobility of the elements. On the other hand, the benchmarks of “DPS” and “Rand” cases demonstrate a decreasing trend with increasing region size, while the “Conven” case has little change as the region size varies. Compared with the benchmarks, the proposed scheme and the optimization algorithm achieve superior performance across all region sizes, which is consistent to the previous simulation results.
V Conclusions
This paper proposed a fRIS-aided ISAC system to achieve multi-user communications and multi-target sensing by exploiting the beamforming and the fRIS element position optimization. We formulated a joint communication symbol estimation MSE and sensing beampattern mismatch minimization problem subject to transmit power budget, constant-modulus constraint and the element position requirements of the fRIS. An AM-based algorithm was developed, where each subproblem utilized the techniques of derivatives, SDR, ALM and MM optimization. Simulation results validated that the fRIS can significantly reduce the sensing beampattern mismatch and communication symbol estimation error in the ISAC systems, as compared to the conventional RIS.
Appendix A Derivation of the function $\widetilde{f}_{0}$ in (III-E)
To simplify the objective function (34a), we first expand the two square norms, and it is given as
$\!\!\!\alpha\cdot\left(\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}-2\mathfrak{R}\left\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right)\!+\!(1\!-\!\alpha)\cdot$
$\!\!\!\!\left(\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\!\!-\!\!2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right).$ (49)
It is noteworthy that $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
Observing the first term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is actually unrelated to the variable $\mathbf{p}$ due to the following reason. Given the property of $\boldsymbol{\Theta}$ that $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (10), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$. Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (8), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of element positions. As a result, this term eventually becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
For the second term $-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can also be substituted, yielding $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$. Leveraging the cyclic property of trace, it can be re-expressed as $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
Focusing on the third term, it can be rewritten as
$\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$ (50a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $c_{0}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}\mathbf{A}_{rc}(\mathbf{p})\boldsymbol{\Sigma}_{rc}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})$ (50b)
$\overset{\mathrm{(b)}}{\longrightarrow}$ $c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}\mathbf{a}_{c,k}(\mathbf{p})\mathbf{a}_{c,k}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})$ (50c)
$\overset{\mathrm{(c)}}{\longrightarrow}$ $c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p}).$ (50d)
In operation (a), the expression of $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms for simplicity. In operation (b), it is derived based on the matrix structure in (14a) and (14b). The operation (c) leverages the diagonal property of $\boldsymbol{\Theta}$, where we denote $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$.
As for the last term in (A), it can be transformed as follows:
$-2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}$ (51a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\right\}$ (51b)
$\overset{\mathrm{(b)}}{\longrightarrow}$ $-2\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\},$ (51c)
where (a) is the substitution operation, while (b) utilizes the cyclic property of trace and the diagonal property of $\boldsymbol{\Theta}$. We also denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
By integrating all the simplified terms, the objective function can be expressed as (III-E). This ends the derivation.
Appendix B Derivation of the function $\widetilde{f}_{1}$ in (III-E)
For the function (III-E), it can be further simplified as follows if one considers the position of $n$-th element position $\mathbf{p}_{n}$ while fixing others. First, examining the first term $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which can be expressed as
$\sum_{n=1}^{N}-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (52a)
$\overset{\mathrm{(a)}}{\longrightarrow}$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (52b)
[…]
In particular, $[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$. Additionally, $\nu_{n}=2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is a variable-independent value. The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, which is given as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$. Moreover, in the operation (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
For the second term, one can rewrite it in element-wise manner, i.e.,
$(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})$ (53a)
$\longrightarrow$ $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{j=1}^{N}\zeta_{k}\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,j}\!\!\!\cdot\left[\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\right]_{i}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\cdot\left[\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\right]_{j}$ (53b)
$\longrightarrow$ $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{j=1}^{N}\zeta_{k}\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,j}\!\!\!$
$\!\!\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}\!\cdot\!e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,j}(\mathbf{p}_{j})-d_{c,k,j}(\mathbf{p}_{j})\right)},$ (53c)
where $d_{c,k,n}(\mathbf{p}_{n})=p_{x,n}\sin(\phi_{c,k})\cos(\psi_{c,k})+p_{y,n}\sin(\psi_{c,k})$. Apparently, when $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$. On the other hand, it can be found that the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering $n$-th element (i.e. $i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (53c) can be reduced to
$(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\sum_{i=1,i\neq n}^{N}2\mathfrak{R}\left\{\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,n}\right.\!\!\!$
$\!\!\left.\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}\cdot e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}.$ (54)
Notably, when fixing the positions of other elements, the term $[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$. We denote the phase of $\widetilde{c}_{i,n,k}$ as $\widetilde{\xi}_{i,n,k}=\angle\widetilde{c}_{i,n,k}$ while $\widetilde{c}_{i,n,k}$ has an amplitude of $1$. As a result, (B) can further be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!2\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (55a)
$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (55b)
where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$ for simplicity.
Focusing on the last term, it can be transformed as follows:
$-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\}$ (56a)
$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (56b)
$\longrightarrow$ $\sum_{k=1}^{K}-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (56c)
$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (56d)
where $s_{n,k}=[\mathbf{S}]_{n,k}$ with the amplitude of $\overline{\rho}_{n,k}=|s_{n,k}|$ and the phase of $\overline{\xi}_{n,k}=\angle s_{n,k}$. We also denote $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$ for brevity.
This completes the derivation.
Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
Given that $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the sub-functions $\widetilde{f}_{1,1}$, $\widetilde{f}_{1,2}$, $\widetilde{f}_{1,3}$ shown in (III-E), the derivative of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is equal to the respective derivatives of the sub-functions.
For the first sub-function $\widetilde{f}_{1,1}$, the first-order derivative with respect to $p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear. As a result, the can be handled by the chain rule of derivative. First, we derive the the first-order derivatives of $\widetilde{f}_{1,1}$:
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,x}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57a)
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,y}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57b)
where $\kappa_{r,x}=\sin(\phi_{r})\cos(\psi_{r})$ and $\kappa_{r,y}=\sin(\psi_{r})$. Based on (57), the second-order derivatives can be derived as
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58a)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\!$ $=\!\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}$
$=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}\kappa_{r,y}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58b)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,y}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right).$ (58c)
For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can also be derived as
$\!\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59a)
$\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59b)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59c)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}$
$\!=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59d)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59e)
where $\Delta\widetilde{\kappa}_{x,k}=\kappa_{r,x}-\kappa_{c,x,k}$ and $\kappa_{c,x,k}=\sin(\phi_{c,k})\cos(\psi_{c,k})$, while $\Delta\widetilde{\kappa}_{y,k}=\kappa_{r,y}-\kappa_{c,y,k}$ and $\kappa_{c,y,k}=\sin(\psi_{c,k})$.
Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given as
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60a)
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60b)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60c)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}$
$=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60d)
$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right).$ (60e)
By substituting the results of the first-order and second-order derivatives to (39) and (40), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained.
This finishes the derivations.

### v3

@@
 Abstract
-A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communications (ISAC) system is proposed to enhance multi-target sensing and multi-user communication.
-Unlike the conventional RIS, the fRIS incorporates movable elements whose positions can be flexibly adjusted to provide extra spatial degrees of freedom.
-In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and communication symbol estimation error by optimizing the symbol estimator, transmit beamformer, fRIS phase shifts, and element positions.
-To solve this problem, an algorithm based on alternating minimization is devised, where subproblems are solved leveraging augmented Lagrangian method, quadratic programming, semidefinite-relaxation, and majorization-minimization techniques.
-A key challenge exists that the fRIS element positions affect both the incident and reflective channels, leading to the high-order composite functions regarding the positions.
-As a remedy, it is proved that the high-order terms can be transformed to linear and linear-difference forms using the characteristics of fRIS and structural channels, which facilitates the position optimization.
-Numerical results validate the effectiveness of the proposed scheme as compared to the conventional RIS-aided ISAC systems and other benchmarks.
-Index Terms:
-Alternating minimization, fluid reconfigurable intelligent surface, integrated sensing and communications, mismatch and estimation error
+A fluid reconfigurable intelligent surface (fRIS)-aided integrated sensing and communication (ISAC) system is proposed to enhance multi-target sensing and multi-user communication.
+Unlike the conventional RIS, the fRIS employs movable elements with adjustable positions, offering additional spatial degrees of freedom.
+In this system, a joint optimization problem is formulated to minimize sensing beampattern mismatch and symbol estimation error.
+An algorithm based on alternating minimization is devised to handle the resultant non-convex problem, where the subproblems are solved via augmented Lagrangian method, quadratic programming, semidefinite relaxation, and majorization-minimization.
+A key challenge is that the element positions affect both incident and reflective channels, leading to the high-order composite objective functions.
+As a remedy, the high-order terms are transformed into linear and linear-difference forms by exploiting the structural characteristics of fRIS and the channels.
+Numerical results demonstrate the superiority of the proposed scheme over conventional RIS-aided ISAC and other benchmarks.
 I Introduction
-In the era of sixth-generation (6G) wireless communications, integrated sensing and communications (ISAC) emerged as a promising paradigm [1], addressing spectrum scarcity and reducing resource consumption.
-The ISAC aims to incorporate the sensing systems and the communication systems at different levels.
-Herein, the two systems shared the same spectrum and coexisted with minimal mutual interference [2] using null space projection [3], transceiver design [4], etc.
-To further mitigate interference, collaborations between the two systems were achieved by sharing critical information    [5], [6].
-Additionally, considering the similarities in hardware and signal processing, the recent ISAC advancement focuses on designing unified platforms that can perform the dual functions [7].
-For example, in [8], the transmit beamformer was designed by combining weighted radar waveforms and communication symbols.
-Besides, a wideband scenario about joint design was considered in [9], where mutual information was utilized as a unified metric to design the ISAC waveform.
-Meanwhile, reconfigurable intelligent surface (RIS) garnered great attention in the 6G innovation.
-Specifically, the RIS consists of numerous reflective elements, each of which is able to induce phase shifts to the incident signals.
-This enables the RIS to dynamically adapt the channel propagation environment by designing the phase shifts [10].
-Leveraging this characteristic, work [11] demonstrated significant improvement in communication sum-rate in RIS-aided multi-user scenarios.
-Apart from this, it was also shown that the RIS can suppress the undesired signals, such as eavesdropping  [12] and jamming [13].
-Additionally, the RIS was applied in other diverse scenarios, including radar detection [14], near-field communications [15], and wireless power transfer and communications [16].
-To harness the combined benefits of ISAC and RIS, extensive research was conducted on RIS-aided ISAC systems [17].
-In [18], authors investigated an echo signal-to-noise ratio (SNR) maximization problem in a RIS-aided ISAC system, which serves a communication user and tracks a target through joint design of the transmit precoder and RIS beamformer.
-Building on this, [19] extended this framework to multi-user and multi-target scenarios under practical constraints.
-Furthermore, authors in [20] and [21] considered the energy efficiency maximization for the RIS-aided ISAC systems without compromising the communication and sensing performance.
-Additionally, researchers examined various other applications of the RIS-aided ISAC systems, such as simultaneously transmitting and reflecting surface [22], unsupervised learning [23], and extended reality [24].
-While the RIS-aided ISAC systems offer significant enhancements for 6G communications, there remains room for the improvement in spatial degrees of freedom (DoFs) of the systems and the reduction in system complexity.
-Particularly, the antennas of the conventional arrays are fixed after assembly, which restricts the spatial diversity [25].
-On the other hand, to achieve excellent performance, a massive number of elements are required, which increases manufacturing costs and exacerbates high-dimensional optimization complexity.
-To address these limitations, movable antenna, also known as fluid antenna, becomes a promising candidate.
-Enabled by the advancement of software-controllable fluidic metamaterials, the movable antenna can dynamically adjust its position in real time, offering additional spatial DoFs  [26], [27].
-Moreover, the enhanced performance brought by the movable antenna allows comparable gains with fewer elements, significantly reducing the computational complexity caused by high-dimensional optimization.
-The potential of the fluid antenna system was extensively studied.
-For instance, researchers demonstrated that it can promote the capacity improvement in point-to-point communication systems [25] and enhance downlink sum-rate in multi-user communication scenarios [28].
-Besides, it was applied in ISAC systems to jointly optimize transceiver designs and antenna positions, enhancing communication and sensing performance [29], [30], [31].
-Meanwhile, efforts to integrate the RIS with the fluid antenna system was made, where authors in [32] and [33] analyzed the outage probability bounds and asymptotic distributions for fluid antenna-aided RIS systems.
-Nevertheless, these works still rely on fixed RIS elements and do not fully exploit the spatial DoFs of RIS elements.
-By contrast, [34] introduced movable RIS elements with flexible positioning and studied the elimination of phase distribution offset in the system.
-However, this work was limited to a simple single-antenna and single-user communication setup, while did not consider the system optimization problem.
-This leaves room for further exploration in the potential of the movable RIS elements in ISAC systems.
-Against the aforementioned background, we are motivated to study an ISAC system aided by a fluid RIS (fRIS).
-In particular, fRIS, named after the fluid antenna system, implies that the elements of RIS are moveable.
-In this system, beams are formed at the fRIS to serve the communication users and illuminate the targets, where the sensing beampattern mismatch and communication decoding error are jointly minimized.
-This problem is solved by an alternating minimization (AM)-based algorithm, which incorporates augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite-relaxation (SDR), and majorization-minimization (MM) technique.
-The main contributions of this work are summarized as:
-  • As the movable elements can provide additional spatial DoFs compared to the conventional RIS, we propose a novel fRIS-aided ISAC system to improve the communication and sensing performance.
-Specifically, the ISAC transmit signal, fRIS beampattern for target illumination, communication symbol decoding and fRIS-related channels are characterized, and a joint sensing beampattern mismatch and communication decoding error minimization problem is formulated.
-  • To tackle the resultant optimization problem, we devise an algorithm based on AM.
-First, we reformulate the original high-order problem into a more manageable form.
-Then, the problem is divided into several subproblems and dedicated optimization approaches are established for each subproblem.
-Eventually, the convergence and the computational complexity of the proposed algorithm are analyzed.
-  • In the subproblem of optimizing the fRIS element positions, the optimization process is challenging since the changing RIS elements simultaneously affects the incident channel, the reflective channels.
-To address this, we leverage the characteristics of the fRIS and the structural channels to reduce the fourth-order terms of the element position to the linear and linear difference terms, so that the objective function of the subproblem can be simplified.
-This facilitates the optimization process.
-  • Numerical results are demonstrated to validate the effectiveness of the proposed model.
-The results reveal a significant performance gain achieved by employing the fRIS in the ISAC system and the proposed scheme outperforms the benchmarks.
-Moreover, by employing the fRIS, the system performance loss caused by the reduction of the element number can be complemented, which alleviates the curse of dimensionality in the RIS optimization.
+In the era of sixth-generation wireless communications, integrated sensing and communications (ISAC) have emerged as a promising paradigm to alleviate resource scarcity [7].
+In ISAC, the communication system and the sensing system aim to share the same spectrum [15], where various schemes have been developed, such as null-space projection [12] and transceiver design [30].
+Furthermore, the works [36, 1] achieved collaborations of the two systems by sharing critical information.
+Additionally, given the hardware and signal processing similarities between sensing and communication, recent ISAC research focused on unified platforms capable of both functions [9].
+For instance, the work [19] designed the transmit beamformer by combining radar waveforms and communication symbols, while the work [40] addressed wideband ISAC using a unified waveform design metric.
+In [25], a latency-aware resource allocation scheme was developed for multiple-input and multiple-output ISAC systems.
+In addition, recent works explored parameter estimation in ISAC systems.
+For example, the work [5] proposed a hybrid fusion framework, where the fused signal was separated for the estimation of the target directions.
+Meanwhile, reconfigurable intelligent surfaces (RIS) have attracted great attention.
+Comprised of many reflective elements, RIS is capable of inducing phase shifts on incident signals and dynamically reconfiguring wireless channels [24].
+This capability enables substantial communication sum-rate gains in multi-user scenarios [35], and facilitates suppression of undesired signals such as eavesdropping [8] and jamming [33].
+Besides, the work [26] showed that RIS can construct favorable channels to support devices with ultra-reliable and low-latency communications.
+In addition, RIS can be applied to various areas, such as radar detection [29], near-field communications [44], and wireless power transfer and communication [45].
+To harness the combined benefits of ISAC and RIS, extensive studies have been conducted on RIS-aided ISAC systems [18].
+In [10], the authors maximized the echo signal-to-noise ratio (SNR) by jointly designing the transmitter and RIS beamformer in a single-user and single-target scenario.
+This framework was extended in [17] to support multi-user and multi-target systems.
+Furthermore, the work [2] addressed the issue of high dynamic range caused by strong interference in RIS-aided ISAC systems, where a joint active and passive beamforming was designed to reduce the dynamic range and mitigate the interference.
+To reduce energy consumption, energy efficiency optimization was explored in RIS-aided ISAC systems [55, 48].
+Additionally, different applications were investigated, including simultaneously transmitting and reflecting surfaces [39], unsupervised learning [47], and extended reality [22].
+In the above works, there remains potential to improve spatial degrees of freedom (DoFs) and reduce system complexity.
+Particularly, the conventional arrays have fixed antennas after assembly, which restricts the spatial diversity [23].
+Furthermore, achieving high performance often requires a large number of elements, leading to increased manufacturing costs and optimization burdens.
+To overcome these issues, the fluid antenna (FA), also known as movable antenna, has emerged as a candidate.
+Enabled by fluidic metamaterials, FA can dynamically adjust the antenna positions, offering additional spatial DoFs [41].
+Moreover, the improved performance enables comparable gains with fewer elements, thereby reducing the computational load of large-scale optimization.
+The potential of FA was extensively explored, such as point-to-point capacity [23], communication sum-rate [27], and ISAC performance improvement [56, 58, 51].
+To adaptively control the propagation environment, some efforts incorporated RIS in FA systems [31, 46].
+However, they still rely on fixed RIS elements.
+Considering that RIS has more elements, fluid RIS (fRIS) or movable-element RIS was studied [32, 43, 6, 57, 53, 54, 20] to exploit more DoFs, where the elements in RIS are movable.
+Compared to the RIS-aided FAS system, fRIS has several advantages.
+First, fRIS can bring more spatial DoFs, as the RIS typically contains more elements than the BS antennas.
+Second, RIS elements are passive and free from the RF chain constraints.
+Third, from a deployment perspective, fRISs are typically more ubiquitous than FAS-BSs in large-scale networks, effectively improving system performance.
+In the fRIS systems, several works [32, 43, 6] explored pixel-based hardware architectures that emulate virtual element movement through densely distributed port switching.
+In contrast, other works [57, 53, 54, 20] employed flexible cables, motors, and sliding tracks to reposition the elements for communication enhancement.
+While these implementations inevitably add hardware complexity, the fRIS can achieve improved performance with fewer elements due to the additional DoFs in a new dimension [^1] The fRIS introduces new spatial DoFs by allowing controlled displacement of each reflecting element as compared to the traditional RIS.
+These position-based DoFs are fundamentally distinct from phase-shift control, as they reshape the propagation geometry, offering a complementary optimization domain.
+.
+Additionally, enabling technologies are actively studied and continuously refined, making the hardware complexity manageable.
+This offers a new direction for RIS evolution.
+However, these works concerning the fRIS mainly focus on basic communication settings and lack ISAC modeling and optimization, motivating further study in this direction.
+Against the above background, we investigate a fRIS-aided ISAC system, where the fRIS forms beams to simultaneously serve communication users and illuminate sensing targets.
+A joint optimization problem is formulated to minimize both sensing beampattern mismatch and communication decoding error.
+Then, the resultant problem is addressed using the alternating minimization (AM)-based algorithm incorporating augmented Lagrangian method (ALM), quadratic programming (quadprog), semidefinite relaxation (SDR), and majorization-minimization (MM).
+The main contributions are summarized as follows:
+  • Leveraging the additional spatial DoFs introduced by movable elements, we propose a novel fRIS-aided ISAC system to enhance sensing and communication.
+Specifically, we characterize the transmit signal, fRIS beampattern, symbol decoding, and fRIS-related channels.
+Besides, we formulate a joint minimization problem for sensing mismatch and decoding error.
+  • To solve the resultant problem, we devise an AM-based algorithm.
+First, the original high-order formulation is transformed into a more tractable form.
+Then, the problem is decomposed into subproblems, which are tackled using dedicated optimization techniques.
+Furthermore, we analyze the convergence and computational complexity of the proposed algorithm.
+  • The subproblem of optimizing fRIS element positions is particularly challenging, as their movement affects both incident and reflective channels.
+To address this issue, we exploit the structural properties of fRIS and channels to reduce the fourth-order terms to linear and linear-difference forms, simplifying the objective and facilitating the optimization.
+  • Numerical results validate the effectiveness of the proposed model, showing notable performance gains over existing benchmarks.
+Moreover, fRIS compensates for performance degradation caused by reducing the number of RIS elements, alleviating the high-dimensional complexity in RIS optimization.
 This paper is organized as follows.
-Section II introduces the model architecture of the fRIS-aided ISAC system, while formulates the symbol error and beampattern mismatch minimization problem.
-In section III, the AM-based algorithm for the resultant problem and the analysis of the proposed algorithm are presented in details.
-Section IV demonstrates the numerical results.
-Finally, conclusions are drawn in Section V.
-[…]
+Section II presents the architecture of the fRIS-aided ISAC system and formulates the joint minimization problem.
+Section III details the proposed AM-based algorithm, while analyzes its convergence and complexity.
+Numerical results are shown in Section IV, and conclusions are drawn in Section V.
+Notation: Bold lowercase and uppercase letters denote vectors and matrices, respectively.
+In addition, $(\cdot)^{\mathrm{H}}$, $(\cdot)^{\mathrm{T}}$, and $(\cdot)^{*}$ signify conjugate transpose, transpose, and conjugate, respectively.
+$\mathrm{tr}(\cdot)$ is the trace of a matrix.
+$\mathrm{diag}(\cdot)$ extracts diagonal entries into a vector, while $\mathrm{Diag}(\cdot)$ creates a diagonal matrix from a vector.
+$\mathrm{vec}(\mathbf{A})$ vectorizes matrix $\mathbf{A}$.
+$|\cdot|$ and $\|\cdot\|$ denote absolute value and norm, respectively.
+$\mathfrak{R}[x]$ and $\angle x$ are the real part and phase of a complex scalar $x$, respectively.
 II System Model and Problem Formulation
@@
 Figure 1: A scenario of a fRIS-aided ISAC system that performs multi-target sensing and multi-user communication.
-As depicted in Fig.1, we consider an ISAC system aided by a fRIS.
-Specifically, the ISAC base station (BS), equipped with a $M$-antenna uniform linear array (ULA), aims to simultaneously serve $K$ single-antenna communication users and sense $T$ point targets.
-Without loss of generalization, it is considered that the direct sensing and communication paths are blocked so that the fRIS is deployed to establish the virtual links and sensing paths.
-The fRIS has $N$ movable elements, where all the elements can move freely within a planar region $\mathcal{A}$ with the shape of $A\times A$.
-For $n$-th fRIS element, its position is denoted by its Cartesian coordinate $\mathbf{p}_{n}=\left[p_{x,n},p_{y,n}\right]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$.
-As a result, the positions of all the elements can be given compactly as $\mathbf{p}=\left[\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{N}\right]\in\mathcal{A}^{2\times N}$.
-In this system, we denote the data symbols for the users as $\mathbf{s}_{c}\in\mathbb{C}^{K\times 1}$, which is precoded by a beamformer $\mathbf{W}\in\mathbb{C}^{M\times K}$ in the BS.
-This yields the precoded digital signal as
-$\mathbf{x}=\mathbf{W}\mathbf{s}_{c}\in\mathbb{C}^{M\times 1}.$ (1)
-The symbol $\mathbf{s}_{c}$ is considered statistically independent and auto-correlated, namely $\mathbb{E}[\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}]=\mathbf{I}_{K\times K}$.
-After converting the radio frequency domain, the signal $\mathbf{x}$ is emitted from the BS and propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS.
-The phase of the incident signal in the fRIS is altered adaptively by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\left\{\theta_{n},\forall n=1,\dots,N\right\}$ represents the phase shift coefficient for $n$-th RIS element.
-As a result, the reflective signal at the fRIS can be given as
+As illustrated in Fig.
+1, we consider an ISAC system assisted by a fRIS.
+The ISAC base station (BS), equipped with a uniform linear array (ULA) of MM antennas, simultaneously serves $K$ single-antenna communication users and senses $T$ point targets.
+Without loss of generality, we assume that direct links for both sensing and communication are blocked and thus a fRIS is deployed to establish virtual communication and sensing paths.
+The fRIS is comprised of $N$ movable elements, where each element can reposition freely within a planar region $\mathcal{A}$ of size $A\times A$.
+The position of the $n$-th element is denoted as $\mathbf{p}_{n}=[p_{x,n},p_{y,n}]^{\mathrm{T}}\in\mathcal{A}^{2\times 1}$, where all the element positions are given by $\mathbf{p}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{N}]\in\mathcal{A}^{2\times N}$.
+In this system, the signal $\mathbf{x}\in\mathbb{C}^{M}$ is transmitted from the BS, which acts as dual-functional waveform for both sensing and communication.
+The transmit signal propagates through the channel $\mathbf{G}\in\mathbb{C}^{N\times M}$ to the fRIS.
+In the fRIS, the phase of the incident signal $\mathbf{G}\mathbf{x}$ is altered by an adjustable phase shift matrix $\mathbf{\Theta}=\mathrm{Diag}\left(\theta_{1},\theta_{2},\cdots,\theta_{N}\right)$, in which $\theta_{n}$ represents the phase shift coefficient for the $n$-th fRIS element.
+As a result, the reflective signal at the fRIS is given by
 $\mathbf{v}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}.$ (1)
-Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and propagate through the reflecting channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ to serve the communication users.
+Subsequently, the reflective signal at the fRIS ought to form multiple beams to illuminate the targets and serve the communication users via the reflecting channel $\mathbf{H}_{rc}$, where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ represents the overall downlink reflective channel while $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$ is the reflective channel from the fRIS to the $k$-th user.
 II-A1 Sensing Model
-As we know, the illumination power to the targets is direction related to the echo SNR, where more target information can obtained if the echo SNR is high.
-As a consequence, beams should be formed at the fRIS towards the directions of the targets to illuminate the targets.
-To this end, one can design the correlation matrix of the reflective signal $\mathbf{v}$ to manipulate the beampattern of the reflective signal.
-Specifically, the relationship between the beampattern and the correlation matrix of the reflective signal can be expressed as
+In the fRIS, beams should be formed towards the directions of the targets for target illuminations.
+To this end, we can design $\mathbf{v}$ to manipulate the beampattern of the reflective signal.
+Specifically, the beampattern of $\mathbf{v}$ is expressed as
 $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ $=\mathbf{a}^{\mathrm{H}}(\phi,\psi,\mathbf{p})\mathbf{R}_{s}\mathbf{a}(\phi,\psi,\mathbf{p}),$ (2)
-in which $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$ is the covariance matrix of signal $\mathbf{v}$.
-The symbol $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS, which is not only dependent on azimuth angle $\phi$ and elevate angle $\psi$, but also the positions of the movable elements $\mathbf{p}$.
-For conciseness, we leave the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in the subsection II-A3.
-For beam design, a common metric is the beampattern similarity, where the mismatch between the practical beampattern $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and the ideal pattern $\mathcal{P}_{d}(\phi,\psi)$ are measured to evaluate the target sensing performance.
-Specifically, the beampattern mismatch can be written as
+which indicates the illumination power of the fRIS at the azimuth angle $\phi$ and the elevation angle $\psi$.
+In (2), we define $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}=\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}$, while the symbol of $\mathbf{a}(\phi,\psi,\mathbf{p})$ represents the steering vector of the fRIS.
+For conciseness, we present the detailed expression of $\mathbf{a}(\phi,\psi,\mathbf{p})$ in Subsection II-A3.
+For beam design, a common metric is the beampattern similarity, which evaluates the mismatch between $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ and an ideal pattern $\mathcal{P}_{d}(\phi,\psi)$.
+The mismatch is quantified as
 $\varepsilon_{r}=\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2},$ (3)
-in which $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote the grid number of the azimuth angles and elevate angles, respectively.
-When $\varepsilon_{r}$ is minimized, the practical beampattern aligns with the ideal beampattern, such that the reflective signal can effectively illuminate the targets.
+where $\beta$ is a scaling factor, while $I_{a}$ and $I_{e}$ denote azimuth and elevation angle sample numbers, respectively.
+Minimizing $\varepsilon_{r}$ encourages $\mathcal{P}_{s}(\phi,\psi,\mathbf{p},\mathbf{R}_{s})$ to closely match $\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})$, enhancing the target illumination performance.
 II-A2 Communication Model
-Apart from illuminating the targets, the reflective signal $\mathbf{v}$ also propagates to the users for communication through the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}\in\mathbb{C}^{1\times N}$.
-The signal received at user $k$ can be given as
-$y_{c,k}=\mathbf{h}_{rc,k}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+n_{c,k},~{}~{}\forall k=1,\cdots,K,$ (5)
-where $n_{c,k}$ represents the additive Gaussian white noise at the receiver of user $k$ with zero mean and $\sigma_{0}^{2}$ variance.
-To estimate the transmitted symbols from the received signal, an estimation factor $\omega$ is introduced, i.e.,
-$\hat{s}_{c,k}=\omega y_{c,k}.$ (6)
-Given the estimated communication symbols, a direct performance indicator for evaluating the communication performance is the mean square error (MSE) between the estimated symbol and the transmitted symbol.
-It can be formulated as
-$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$ (7a)
-=$\overset{(a)}{=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (7b)
-where $\mathbf{H}_{rc}=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,k}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}\in\mathbb{C}^{N\times K}$ incorporates all the reflective channels.
-The operation $(a)$ substitutes (5) and (6) to (7a) and leverages the statistical independence and the auto-correlation characteristics of $\mathbf{s}_{c}$.
+After $\mathbf{v}$ propagates to the communication users, the overall received signal $\mathbf{y}_{c}\in\mathbb{C}^{K}$ at the users is given by
+$\mathbf{y}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+\mathbf{n}_{c},$ (4)
+where $\mathbf{n}_{c}\in\mathbb{C}^{K}$ represents the additive Gaussian white noise matrix, whose entries are independent and identically distributed with zero mean and $\sigma_{0}^{2}$ variance.
+We denote $y_{c,k}$ as the received signal of the $k$-th user, which is the $k$-th entry of $\mathbf{y}_{c}$.
+Given a desired constellation symbol ${{s}}_{c,k}$ for the $k$-th user, an estimation factor $\omega$ is introduced to recover ${{s}}_{c,k}$ from $y_{c,k}$, which is given by
+$\hat{s}_{c,k}=\omega y_{c,k},$ (5)
+where $\hat{s}_{c,k}$ is the estimated symbol. To evaluate the communication performance, a direct metric is the mean square error (MSE) between the estimated symbol and the transmitted symbol:
+$\varepsilon_{c}$ $=\mathbb{E}[\|\mathbf{s}_{c}-\hat{\mathbf{s}}_{c}\|^{2}]$
+${=}\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2},$ (6)
+where $\mathbf{s}_{c}\in\mathbb{C}^{K}$ and $\hat{\mathbf{s}}_{c}\in\mathbb{C}^{K}$ incorporate all the ${{s}}_{c,k}$ and $\hat{s}_{c,k}$, respectively.
+In (II-A2), the first term can be viewed as the multi-user interference, while the second term is the impact of the noise.
 II-A3 Channel Model
-As stated before, the steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$ as well as the azimuth angle $\phi$ and the elevate angle $\psi$.
-Specifically, the steering vector of the fRIS can be expressed as [25]:
-$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!.$ (8)
-The symbol $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, which gives
+The steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$, the azimuth angle $\phi$, and the elevation angle $\psi$.
+Specifically, the steering vector of the fRIS is expressed as [23]:
+$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!,$ (7)
+where $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, given by
 $d_{n}(\phi,\psi,\mathbf{p}_{n})=p_{x,n}\sin(\phi)\cos(\psi)+p_{y,n}\sin(\psi).$ (8)
-Without loss of generalization, the fRIS is typically deployed in the area where line-of-sight (LoS) paths dominate, so that the signal coverage can be extended.
-Consequently, the fRIS-related channels are modeled as LoS channels, which are determined by the steering vector of the fRIS.
-For the channel $\mathbf{G}$ from the BS to the fRIS, it can be modeled as
+Without loss of generality, the fRIS is considered deployed in the area where line-of-sight (LoS) paths dominate, such that the signal coverage can be extended.
+Consequently, the fRIS-related channels are modeled as LoS channels.
+The channel from the BS to the fRIS is modeled as
 $\mathbf{G}$ $=\sqrt{\zeta_{G}}\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}(\phi_{t}),$ (9)
-in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth direction and $\psi_{r}$ in elevate direction, which is in the form of (8).
-On the other hand, the ULA is equipped in the BS, so that the transmit steering vector $\mathbf{a}_{t}$ in the direction of $\phi_{t}$ can be written as
+in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth and $\psi_{r}$ in elevation.
+On the other hand, as the ULA is equipped in the BS, the transmit steering vector in the direction of $\phi_{t}$ is written as
 $\mathbf{a}_{t}=\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{1}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{m}}&\cdots&e^{\jmath\frac{2\pi}{\lambda}\widetilde{d}_{M-1}}\end{bmatrix}^{\mathrm{T}},$ (10)
 where $\{\widetilde{d}_{m}=\frac{\lambda}{2}m\sin(\phi_{t}),~\forall m=0,\cdots,M-1\}$ and the inter-element spacing is half wavelength.
-Similarly, the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}$ from the fRIS to the user $k$ is modeled as
+Similarly, the channel from the fRIS to the $k$-th user is expressed as
 $\mathbf{h}_{rc,k}^{\mathrm{H}}=\sqrt{\zeta_{k}}\mathbf{a}^{\mathrm{H}}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p}),\forall k=1,\cdots,K,$ (11)
-where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the transmit steering vector of the fRIS to user $k$ from azimuth direction $\phi_{c,k}$ and elevate direction $\psi_{c,k}$, which takes the form of (8).
-The path loss between the fRIS and the user $k$ is denoted as $\sqrt{\zeta_{k}}$.
-Accordingly, the overall reflective channel $\mathbf{H}_{rc}$ can be rewritten as
+where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the steering vector of the fRIS to the $k$-th user from $\phi_{c,k}$ and $\psi_{c,k}$.
+The path loss between the fRIS and the $k$-th user is denoted as $\sqrt{\zeta_{k}}$.
+Accordingly, the overall downlink reflective channel $\mathbf{H}_{rc}$ is rewritten as
 $\mathbf{H}_{rc}$ $=\begin{bmatrix}\mathbf{h}_{rc,1}&\!\!\!\cdots\!\!\!&\mathbf{h}_{rc,K}\end{bmatrix}=\mathbf{A}_{rc}\boldsymbol{\Sigma}_{rc},$ (12)
-where we denote
+where $\mathbf{A}_{rc}$ and $\mathbf{\Sigma}_{rc}$ are defined as
 $\mathbf{A}_{rc}$ $=\begin{bmatrix}\mathbf{a}_{c,1}(\phi_{c,1},\psi_{c,1},\mathbf{p})&\!\!\!\cdots\!\!\!&\mathbf{a}_{c,K}(\phi_{c,K},\psi_{c,K},\mathbf{p})\end{bmatrix}$, (13a)
-$\boldsymbol{\Sigma}_{rc}$ $=\mathrm{Diag}\left(\begin{bmatrix}\sqrt{\zeta_{1}}&\cdots&\sqrt{\zeta_{K}}\end{bmatrix}\right).$ (14b)
+[…]
+Note: In practice, RIS element positions may deviate slightly from their designed values due to hardware imperfections.
+Therefore, deviation $\mathbf{p}=\overline{\mathbf{p}}+\Delta\mathbf{p}$ may introduce phase shifts offset in the steering vector $\mathbf{a}(\mathbf{p})=\boldsymbol{\Pi}(\Delta\mathbf{p})~\overline{\mathbf{a}}(\overline{\mathbf{p}})$.
+It can be analyzed that the effect on communication is negligible due to phase cancellation, while the sensing performance may experience distortion in the beampattern shape, which remains limited under small deviations.
+While the extension to a robust framework regarding position deviations is non-trivial, it is left to future work.
 II-B Problem Formulation
-In order to simultaneously achieve the sensing and the communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (4) and the communication symbol estimation MSE $\varepsilon_{c}$ of (7) is minimized by optimizing the transmit beamformer $\mathbf{W}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ as well as the scaling factor $\beta$ of the sensing beampattern.
-Consequently, the problem can be formulated as
-$\min_{\mathbf{W},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~{}\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
+To simultaneously achieve sensing and communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (3) and the communication symbol estimation MSE $\varepsilon_{c}$ of (II-A2) is minimized by optimizing the transmit waveform $\mathbf{x}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ and the scaling factor $\beta$ of the sensing beampattern.
+Consequently, the problem is formulated as
+$\min_{\mathbf{x},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
 $\!\!+\!(1\!-\!\alpha)\!\left(\!\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\!\right)$ (14a)
-$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (15b)
+$\mathrm{s.t.}$ $~\|\mathbf{x}\|^{2}\leq P_{t},$ (14b)
 $~\omega\in\mathbb{R},$ (14c)
@@
 $~\|\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}\!\geq\!\Delta D,\forall n,n^{\prime}\!\!=\!\!1,\!\cdots\!,N,n\!\neq\!n^{\prime},$ (14f)
-where $\alpha$ is the weighting coefficient indicating the bias of the sensing and the communication.
-The constraint (15b) represents that the transmit power should not exceed $P_{t}$, while in the constraint (15d), the modulus-one phase shift of the fRIS is required.
-The constraints (15e) and (15f) indicate that each RIS element ought to move within the pre-set planar region, and the inter-space among elements should be larger than $\Delta D$.
-It is challenging to solve the problem (15) owing to the two reasons.
-On the one hand, (15d) is a constant modulus constraint, while (15f) takes the non-convex form of “Norm $\geq$ Constant”.
-Besides, the objective function is highly complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to different variables.
+where $\alpha$ is the weighting coefficient, controlling the trade-off between sensing and communication.
+The constraint of (14b) indicates that the transmit power should not exceed $P_{t}$, while (14d) requires the modulus-one phase shifts of the fRIS.
+The constraints of (14e) and (14f) are introduced to model the physical limitations of the fRIS structure.
+Specifically, (14e) captures the physical boundary of the movement area, ensuring that all elements are confined in $\mathcal{A}$.
+Besides, (14f) enforces a minimum distance $\Delta D$ among elements, mitigating the electromagnetic mutual coupling effects and avoiding element collision.
+The problem (14) is non-convex owing to the two reasons.
+On the one hand, (14d) and (14f) are non-convex.
+Besides, the objective function is complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to (w.r.t.) different variables, while the variables are coupled and not jointly convex.
 III Proposed Algorithm
-In this section, an effective AM-based algorithm is developed to address the problem (15) by iteratively updating the variables, and we analyze the convergence and computational complexity of the proposed algorithm.
+In this section, an AM-based algorithm is developed to address (14).
+Given the non-convexity and NP-hardness of the problem, we solve it by iteratively updating the variables.
+Additionally, we analyze the convergence and the computational complexity of the proposed algorithm.
 III-A Reformulation of $\varepsilon_{r}$
-Carefully inspecting $\varepsilon_{r}$ in (4), it can be observed that $\varepsilon_{r}$ is a quadratic function with respect to the covariance matrix $\mathbf{R}_{s}$, while $\mathbf{R}_{s}$ is also a quadratic function with respect to the variables, making $\varepsilon_{r}$ a quartic function with respect to the variables.
-To simplify the objective function in (15a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function with respect to the variables.
-Specifically, we first design an optimum covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only takes the sensing into consideration.
-According to [35], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can given as
+Carefully inspecting $\varepsilon_{r}$ in (3), it is a quartic function of the optimization variables, as it is quadratic w.r.t.
+$\mathbf{R}_{s}$, which itself depends quadratically on the optimization variables.
+To simplify the objective function in (14a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function w.r.t.
+the variables.
+First, we design a reference covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only considers sensing.
+According to [16], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can be given by
 $\min_{\beta,\widetilde{\mathbf{R}}_{s}}$ $~\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}},\widetilde{\mathbf{R}}_{s})\right|^{2}$ (15a)
@@
 $\mathrm{rank}\left(\widetilde{\mathbf{R}}_{s}\right)=1,$ (15d)
-where the objective function requires the designed sensing beampattern to be similar to an ideal one, the constraint (16b) indicates that $\widetilde{\mathbf{R}}_{s}$ is semi-definite, the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$ in the constraint (16c), and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since the ultimately designed $\mathbf{R}_{s}$ should satisfy $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$.
-The problem (16) can be directly solved with SDR approach.
-To guarantee the rank-one constraint of $\widetilde{\mathbf{R}}_{s}$, various methods can be employed, such as Gaussian randomization [36] and penalty-based method [37].
-With the obtained $\widetilde{\mathbf{R}}_{s}$ designed in (16), the beams can be effectively formed towards the targets.
-Considering that the designed $\widetilde{\mathbf{R}}_{s}$ is rank-one, the corresponding reflective signal at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be retrieved by the performing eigenvalue decomposition to $\widetilde{\mathbf{R}}_{s}$.
-This yields
+where the objective function requires the designed sensing beampattern to be similar to an ideal one.
+In the constraint of (15c), the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$, and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$.
+This problem can be directly solved via SDR with Gaussian randomization [21] and penalty-based method [52].
+Remark 1: To handle (15d), SDR followed by Gaussian randomization can be adopted to construct a feasible rank-one solution.
+While a theoretical proof of tightness is generally intractable, this approach is widely adopted and has demonstrated strong empirical performance.
+For the penalty-based method, the rank-one constraint can be enforced via a penalty term $\big(\mathrm{tr}(\widetilde{\mathbf{R}}_{s})-\|\widetilde{\mathbf{R}}_{s}\|_{2}\big)$.
+As proven in [50], a rank-one solution can be obtained when the penalty factor is sufficiently large.
+With the obtained $\widetilde{\mathbf{R}}_{s}$ in (15), the beams can be effectively formed towards the targets.
+Since $\widetilde{\mathbf{R}}_{s}$ is rank-one, a reference signal $\mathbf{s}_{r}$ at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be obtained by the eigenvalue decomposition of $\widetilde{\mathbf{R}}_{s}$, yielding
 $\widetilde{\mathbf{R}}_{s}=\mathbf{s}_{r}\mathbf{s}_{r}^{\mathrm{H}},$ (16)
-where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$, while $\lambda_{s}$ and $\mathbf{u}_{s}$ represent the maximum eigenvalue and the corresponding eigenvector of $\widetilde{\mathbf{R}}_{s}$, respectively.
-The factorized signal $\mathbf{s}_{r}$ represents an optimum and achievable reflective signal whose correlation matrix is $\widetilde{\mathbf{R}}_{s}$ producing directive beams towards the targets.
-In other words, $\mathbf{s}_{r}$ can be viewed as a desired reflective signal for sensing the targets.
-Consequently, the sensing metric is reformulated as MSE between the practical reflective signal $\mathbf{v}$ at the fRIS and the desired reflective signal $\mathbf{s}_{r}$, i.e.
+where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$ with $\lambda_{s}$ and $\mathbf{u}_{s}$ being the maximum eigenvalue and the associated eigenvector, respectively.
+Here, $\mathbf{s}_{r}$ represents a desired and achievable reflective signal, which produces directive beams towards the targets.
+Consequently, the sensing metric is reformulated as the mismatching between $\mathbf{v}$ and $\mathbf{s}_{r}$, i.e.,
 $\widetilde{\varepsilon}_{r}=\|\mathbf{s}_{r}-\mathbf{v}\|^{2}=\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (17)
-It can be observed that the reformulated sensing metric $\widetilde{\varepsilon}_{r}$ becomes quadratic with respect to all the variables, which simplifies the subsequent optimization.
-Ultimately, the problem (15) can be reformulated as
-$\min_{\mathbf{W},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~{}\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
+It is observed that $\widetilde{\varepsilon}_{r}$ becomes quadratic w.r.t. all variables, which simplifies the subsequent optimization.
+Remark 2: The reformulation approximation may incur error from two sources: (i) the rank-one recovery of $\widetilde{\mathbf{R}}_{s}$ via decomposition, which is mainly from the potential suboptimality of Gaussian randomization, and (ii) the imperfect realization of $\mathbf{s}_{r}$ via $\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$ under the joint ISAC design.
+The dominant error arises from the mismatch $\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}$, which is minimized.
+Although theoretical bounds are difficult to obtain, the approximation error remains trivial in practice.
+Ultimately, (14) is reformulated as
+$\min_{\mathbf{x},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
 $+(1\!-\!\alpha)\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!K\omega^{2}\sigma_{0}^{2}\right)$ (18a)
 […]
-However, the problem is still challenging owing to the non-convex constraints and objective function.
-In the subsequence, the problem (19) is solved by iteratively updating each variable with others being fixed.
-It is worth mentioning that $\mathbf{s}_{r}$ may change after the update of $\mathbf{p}$, but according to the Lyapunov stability [38], $\mathbf{s}_{r}$ is asymptotically stable over iterations due to the bounded $\mathbf{p}$ and the finite resources.
+Subsequently, we solve (18) by iteratively updating the variables.
 III-B Optimization of Communication Symbol Estimator $\omega$
-When $\mathbf{x}$, $\mathbf{p}$ and $\boldsymbol{\Theta}$ are fixed, the objective function in the problem (19) is a quadratic function with respect to $\omega$.
-After ignoring the terms that are unrelated to $\omega$, the problem can be simplified as
+Concerning $\omega$, (18a) is a quadratic function. After ignoring the terms that are unrelated to $\omega$, (18) is simplified as
 $\min_{\omega\in\mathbb{R}}$ $~~\varepsilon_{c}=\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}+K\omega^{2}\sigma_{0}^{2}.$ (19)
-The optimal $\omega^{\star}$ can therefore be derived by letting the derivative of $\varepsilon_{c}$ be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$.
-This yields a closed-form optimum:
+The optimal $\omega^{\star}$ can be derived by setting the derivative of $\varepsilon_{c}$ to be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$, yielding the closed-form solution:
 $\omega^{\star}=\frac{\mathfrak{R}\{\mathbf{s}^{\mathrm{H}}_{c}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}}{\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}+K\sigma_{0}^{2}}.$ (20)
 III-C Optimization of fRIS Phase Shift $\boldsymbol{\Theta}$
-Given $\mathbf{W}$, $\mathbf{p}$ and $\omega$, the problem (19) can be reduced to a constant-modulus constrained quadratic minimization problem, which is given as
+Given $\mathbf{x}$, $\mathbf{p}$, and $\omega$, (18) can be reduced to a modulus-one constrained quadratic minimization problem, given by
 $\min_{\boldsymbol{\Theta}}$ $~\alpha\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\!+\!(1\!-\!\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}.$ (21a)
 $\mathrm{s.t.}$ $~|\boldsymbol{\Theta}|_{n,n}=1,~~\forall~n=1,\cdots,N.$ (21b)
-The challenge of this problem mainly stems from the non-convex constant-modulus constraint.
-To address this task, one can first let $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [21] to transform the problem (22) to a compact form:
+To address (21), we first define $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [48] to transform (21) to a compact form:
 $\min_{\boldsymbol{\theta}}$ $~\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{A}}_{1}\boldsymbol{\theta}-\boldsymbol{\theta}^{\mathrm{H}}{\mathbf{b}}_{1}-{\mathbf{b}}_{1}^{\mathrm{H}}\boldsymbol{\theta}$ (22a)
 $\mathrm{s.t.}$ $~|\boldsymbol{\theta}|_{n}=1,\forall n=1,\cdots,N.$ (22b)
-where
-${\mathbf{A}}_{1}$ $\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]$
-$~{}~{}+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right],$ (24a)
-${\mathbf{b}}_{1}$ $=\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})+\omega(1-\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}).$ (24b)
-The nonconvex constant-modulus constraint (23b) remains challenging, where various approaches have been developed in the wide literature to effectively handle such a constraint, such as SDR [39] and manifold optimization [40].
-Therefore, the details will not be elaborated upon here.
-III-D Optimization of Transmit Beamformer $\mathbf{W}$
-In this subsection, the optimization of the transmit beamformer $\mathbf{W}$ will be discussed. Viewing other variables as constant, the problem (19) can be reduced to
-$\min_{\mathbf{W}}$ $~{}\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)$ (25a)
-$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (25b)
-where we denote $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ as the overall channel from the transmitter to the users.
-For the objective function (25a), the norm operations are expanded and thus it becomes
-$\alpha\!\left(\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\}\right)$
-$+$ $(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\}\right).$ (26)
-By leveraging the cyclic property of trace, the relationship of $\mathrm{tr}(\mathbf{X}^{\mathrm{H}}\mathbf{C}_{1}\mathbf{X}\mathbf{C}_{2})=\mathrm{vec}^{\mathrm{H}}(\mathbf{X})(\mathbf{C}_{2}^{\mathrm{T}}\otimes\mathbf{C}_{1})\mathrm{vec}(\mathbf{X})$ and $\mathrm{tr}(\mathbf{C}^{\mathrm{H}}\mathbf{X})=\mathrm{vec}^{\mathrm{H}}(\mathbf{C})\mathrm{vec}(\mathbf{X})$, the objective function can further be simplified as
-$\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\},$ (27)
-in which
-$\mathbf{w}$ $=\mathrm{vec}(\mathbf{W}),$ (28a)
-$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{G}^{\mathrm{H}}\mathbf{G}\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right),$ (28b)
-$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\mathbf{s}_{c}^{\mathrm{H}}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}\right).$ (28c)
-It is worth noting that in (28b), $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ has been simplified as $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ since $\boldsymbol{\Theta}$ is a modulus-one diagonal matrix and $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ always holds.
-After this transformation, it can be seen that the objective (27) is a convex quadratic function.
-As for the constraint (25b), it can also be re-expressed as the inner product of $\mathbf{w}$, that is, $\mathbf{w}^{\mathrm{H}}\mathbf{w}\leq P_{t}$.
-In this context, we consider full power usage in the transmitter, namely $\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t}$.
-As a result, the problem (25) can be written as
-$\min_{\mathbf{w}}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\}$ (29a)
-$\mathrm{s.t.}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t},$ (29b)
-which can be addressed by ALM [41, 42].
-Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part.
-The problem can thus be converted to its real form as
-$\min_{\overline{\mathbf{w}}}$ $~{}\mathring{f}(\overline{\mathbf{w}})=\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{w}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{w}}$ (30a)
-$\mathrm{s.t.}$ $~{}\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}=P_{t},$ (30b)
-where
-$\overline{\mathbf{w}}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{w}\}\\ \mathfrak{I}\{\mathbf{w}\}\end{bmatrix},\overline{\mathbf{b}}_{2}=\begin{bmatrix}\mathfrak{R}\{\mathbf{b}_{2}\}\\ \mathfrak{I}\{\mathbf{b}_{2}\}\end{bmatrix},$ (31a)
-$\overline{\mathbf{A}}_{2}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{A}_{2}\}&-\mathfrak{I}\{\mathbf{A}_{2}\}\\ \mathfrak{I}\{\mathbf{A}_{2}\}&\mathfrak{R}\{\mathbf{A}_{2}\}\end{bmatrix}.$ (31b)
-Accordingly, the augmented Lagrangian function of problem (30) can be written as
-$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{w}})+\mu\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)^{2},$ (32)
-in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. This can be tackled by the following iteration procedure:
-$\overline{\mathbf{w}}^{(t)}$ $=\arg\min_{\overline{\mathbf{w}}}\mathcal{L}(\overline{\mathbf{w}},\mu^{(t-1)}),$ (33a)
-$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{w}}^{\mathrm{T}(t)}\overline{\mathbf{w}}^{(t)}-P_{t}\right).$ (33b)
-Specifically, one can solve the subproblem (33a) by quasi-Newton method [42], which is implemented by the fminunc tool.
-This process is terminated when $\left|\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right|<\epsilon_{0}$ is satisfied.
-It is noted that the penalty factor $\gamma$ gradually increases among the iterations for fast convergence.
-After obtaining $\overline{\mathbf{w}}^{\star}$, it is converted to the complex matrix form and reshaped as $\mathbf{W}^{\star}$.
+where ${\mathbf{b}}_{1}\!=\!\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})\!+\omega(1\!-\!\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}})$ and ${\mathbf{A}}_{1}\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right]$.
+To address (22b), various approaches can be adopted, such as SDR [42] and manifold optimization [49].
+While SDR produces suboptimal solutions due to rank-one relaxation, it ensures a worst-case $\pi/4$-approximation.
+Manifold optimization performs gradient descent over the complex circle manifold, transforming the problem into an unconstrained search.
+Despite inherent non-convexity, both methods achieve high-quality local optima and demonstrate strong empirical performance.
+III-D Optimization of Transmit Waveform $\mathbf{x}$
+In this subsection, the optimization of the transmit waveform $\mathbf{x}$ is discussed. First, (18) w.r.t. $\mathbf{x}$ is simplified as
+$\min_{\mathbf{x}}$ $~\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{x}\|^{2}\right)$ (23a)
+$\mathrm{s.t.}$ $~\|\mathbf{x}\|^{2}\leq P_{t},$ (23b)
+where $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\in\mathbb{C}^{K\times M}$ is denoted as the overall channel from the transmitter to the users.
+For (23a), the squared norm is expanded, and the constant terms are omitted, leading to the equivalent expression as
+$\alpha\!\left(\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}\right)$
+$\!\!\!+(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{x}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{x}\}\right).$ (24)
+The objective function of (III-D) is compactly rewritten as
+$\mathbf{x}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{x}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{x}\}$ (25)
+by denoting
+$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\!\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\!\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right)\!,$ (26a)
+$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\right).$ (26b)
+Notably, when $\boldsymbol{\Theta}$ is an ideal modulus-one diagonal matrix, $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ in (26b) can be reduced to $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ due to $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$.
+After this transformation, it is seen that (25) is a convex quadratic function.
+For (23b), it can be re-expressed as the inner product of $\mathbf{x}$, that is, $\mathbf{x}^{\mathrm{H}}\mathbf{x}\leq P_{t}$.
+In this context, we consider full power usage, namely, $\mathbf{x}^{\mathrm{H}}\mathbf{x}=P_{t}$.
+Therefore, (23) is rewritten as
+$\min_{\mathbf{x}}$ $~\mathbf{x}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{x}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{x}\}~~~~\mathrm{s.t.}~\mathbf{x}^{\mathrm{H}}\mathbf{x}=P_{t},$ (27)
+which can be addressed by ALM [13, 3]. Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part. Thus, (27) is converted to its real form as
+$\min_{\overline{\mathbf{x}}}$ $~\mathring{f}(\overline{\mathbf{x}})=\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{x}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{x}}~~~~\mathrm{s.t.}~\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}=P_{t},$ (28)
+where $\overline{\mathbf{x}}=\left[\mathfrak{R}\{\mathbf{x^{\mathrm{T}}}\},\mathfrak{I}\{\mathbf{x^{\mathrm{T}}}\}\right]^{\mathrm{T}}$, $\overline{\mathbf{b}}_{2}=\left[\mathfrak{R}\{\mathbf{x^{\mathrm{T}}}\},\mathfrak{I}\{\mathbf{x^{\mathrm{T}}}\}\right]^{\mathrm{T}}$ and $\overline{\mathbf{A}}_{2}=\left[\mathfrak{R}\{\mathbf{A}_{2}\},-\mathfrak{I}\{\mathbf{A}_{2}\};\mathfrak{I}\{\mathbf{A}_{2}\},\mathfrak{R}\{\mathbf{A}_{2}\}\right]$.
+Accordingly, the augmented Lagrangian function of (28) can be written as
+$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{x}})+\mu\left(\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right)^{2},$ (29)
+in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. Then, we tackle (29) using the following iteration procedure:
+$\overline{\mathbf{x}}^{(t)}$ $=\arg\min_{\overline{\mathbf{x}}}\mathcal{L}(\overline{\mathbf{x}},\mu^{(t-1)}),$ (30a)
+$\mu^{(t)}$ $=\mu^{(t-1)}+\gamma\left(\overline{\mathbf{x}}^{\mathrm{T}(t)}\overline{\mathbf{x}}^{(t)}-P_{t}\right).$ (30b)
+Specifically, we can solve (30a) via the quasi-Newton method [3], which is implemented by the fminunc tool.
+This process is terminated when $\left|\overline{\mathbf{x}}^{\mathrm{T}}\overline{\mathbf{x}}-P_{t}\right|<\epsilon_{0}$ is satisfied.
 III-E Optimization of fRIS Element Positions $\mathbf{p}$
-In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$ with other variables being fixed.
-First, the problem (19) can be simplified by omitting the unrelated terms and constraints:
+In this subsection, we focus on the optimization of the fRIS element position $\mathbf{p}$. First, (18) is simplified by omitting the unrelated terms and constraints:
 $\min_{\mathbf{p}}$ $~\alpha\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$
 $~+(1-\alpha)\|\mathbf{s}_{c}-\omega\mathbf{H}_{rc}^{\mathrm{H}}(\mathbf{p})\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}(\mathbf{p})\mathbf{x}\|^{2}$ (31a)
-$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},\forall n=1,\cdots,N,$ (34b)
-$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}\geq\Delta D,\forall n,n^{\prime}=1,\cdots,N,n\neq n^{\prime}.$ (34c)
-As mentioned in section II-A3, the channel $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the fRIS element positions.
-It is challenging to solve the problem (34) owing to two main reasons.
-First, the objective function is a complicated composite function concerning the variable $\mathbf{p}$.
-Since the fRIS element positions affect both the incident channel $\mathbf{G}$ and the reflective channel $\mathbf{H}_{rc}$, there appear quadratic and quartic forms simultaneously in the outer function.
-In the interior function, the variable $\mathbf{p}$ is contained in the complex exponential function of an array manifold, shown in (8).
-On the other hand, the constraint (34c) takes the form of “Norm $\geq$ Constant”, rendering it non-convex.
-To explicitly observe $\mathbf{p}$ in (34a), we process the outer function by substituting $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) to (34a) and integrating the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
-Proposition 1. After omitting the constants, the objective function (34a) can be rewritten as
+[…]
+Inspecting (31), $\mathbf{G}$ and $\mathbf{H}_{rc}$ are determined by the variable $\mathbf{p}$, and (31a) is a complicated composite function, where quadratic and quartic forms exist simultaneously in the outer function.
+In the interior function, $\mathbf{p}$ is contained in the complex exponential function, shown in (7).
+To explicitly observe $\mathbf{p}$ in (31a), we substitute $\mathbf{G}$ in (9) and $\mathbf{H}_{rc}$ in (10) to (31a) and integrate the unrelated symbols of $\mathbf{p}$, which is given by the following proposition.
+Proposition 1. After omitting the constants, (31a) is rewritten as
 $\widetilde{f}_{0}=$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\right\}$
@@
 where we denote $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$, $c_{0}=\omega^{2}\zeta_{G}c_{x}$, $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$, $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$, $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
-Proof: Detailed derivations are provided in Appendix A.
-However, the variable in (III-E) is contained in the array manifold, while there still exist $K$ quartic function terms, a quadratic function term and a linear function item with respect to the array manifold.
-Fortunately, it can be proved in the following proposition that the quartic, quadratic and linear items in (III-E) can be reduced to the first-order items and the difference of the first-order items if we only focus on optimizing the position of one element at a time.
-Proposition 2. When considering the position of $n$-th element $\mathbf{p}_{n}$ while fixing others, the objective function $\widetilde{f}_{0}$ in (III-E) can be further simplified as
+Proof: The proof is provided in Appendix A.
+However, $\mathbf{p}$ in (III-E) is contained in the array manifold, while there exist $K$ quartic terms, a quadratic term, and a linear item w.r.t.
+the array manifold.
+Fortunately, it is shown in the following proposition that these items can be reduced to the first-order items and the difference of the first-order items if we focus on optimizing the position of one element at a time.
+Proposition 2. When considering the position of the $n$-th element $\mathbf{p}_{n}$ while fixing others, $\widetilde{f}_{0}$ in (III-E) is simplified as
 $\widetilde{f}_{1}(\mathbf{p}_{n})$ $=\widetilde{f}_{1,1}+\widetilde{f}_{1,2}+\widetilde{f}_{1,3}$
 $=-\nu_{n}\cos\left(\xi_{n}+\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})\right)$
-$~{}~{}+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
-$~{}~{}-\sum_{k=1}^{K}\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (36)
-In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of $n$-th entry of ${}\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant.
-The symbol $d_{r,n}(\mathbf{p}_{n})$ is $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$.
-In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$.
-Meanwhile, we denote $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$.
-In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$ and $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$.
-Proof: Detailed derivations and symbol explanations are shown in Appendix B.
-Observing the objective function $\widetilde{f}_{1}$, it contains cosine functions, which are neither convex nor concave with respect to $\mathbf{p}_{n}$.
-To address this problem, we employ the MM framework, where surrogate functions that locally upper bound the objective function $\widetilde{f}_{1}$ are constructed.
-Specifically, the constructions of the surrogate function is given by the following lemma.
-Lemma 1. [43] Considering a function $\widetilde{f}$ with $\mathbf{x}$ being the variable, its second-order Taylor expansion at the given point $\mathbf{x}_{0}$ can be given as
-$\widetilde{f}(\mathbf{x})\approx\widetilde{f}(\mathbf{x}_{0})$ $+(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})$
-$+\frac{1}{2}(\mathbf{x}-\mathbf{x}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{x}_{0})(\mathbf{x}-\mathbf{x}_{0}),$ (37)
-where $\nabla\widetilde{f}(\mathbf{x})$ is the gradient vector of $\widetilde{f}(\mathbf{x})$, while $\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$ represents the Hessian matrix of the function.
-By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{x}_{0})$, the function $\widetilde{f}(\mathbf{x})$ can be locally upper-bounded by
-$\!\widetilde{f}(\mathbf{x})\leq\widetilde{f}(\mathbf{x}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{x}_{0}))^{\mathrm{T}}(\mathbf{x}-\mathbf{x}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{x}-\mathbf{x}_{0})\right\|^{2}.$ (38)
-In particular, the parameter $\delta$ can be chosen as the Frobenius norm of the Hessian matrix $\nabla^{2}f(\mathbf{x}_{0})$.
-According to the Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function.
-To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined.
-Specifically, the gradient and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ are given by the following theorem.
-Theorem 1.
-According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which can be expressed as
+$~~+\sum_{k=1}^{K}\sum_{i=1,i\neq n}^{N}\widetilde{\nu}\zeta_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right)$
+$~~-\sum_{k=1}^{K}\overline{\nu}~\overline{\rho}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right).$ (33)
+In the first term, we denote $\xi_{n}$ and $\rho_{n}$ as the phase and amplitude of the $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\nu_{n}\!=\!2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is the constant.
+The symbol of $d_{r,n}(\mathbf{p}_{n})$ is defined as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})\!+\!p_{y,n}\sin(\psi_{r})$.
+In the second term, $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\!=\!d_{r,n}(\mathbf{p}_{n})\!-\!d_{c,k,n}(\mathbf{p}_{n})$, while $\widetilde{\xi}_{i,n,k}$ is the phase of a constant $\widetilde{c}_{i,n,k}$.
+Meanwhile, we denote $\widetilde{\nu}=2(1\!-\!\alpha)c_{0}$.
+In the last term, we represent $\overline{\xi}_{n,k}$ as the phase of $s_{n,k}$, while $\overline{\nu}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}$.
+Proof: The proof is provided in Appendix B.
+Remark 3: The derivations in Propositions 1 and 2 rely on structural assumptions of both the fRIS and the channel model.
+Specifically, the fRIS is required to be passive and single-connected, i.e., $|\boldsymbol{\Theta}_{i,i}|=1$.
+Furthermore, the channels are considered as LoS channels, which is reasonable in practical scenarios.
+The RIS is strategically deployed to establish virtual LoS links, for example, to bypass obstacles or extend coverage.
+Observing $\widetilde{f}_{1}$ in (III-E), it contains cosine functions, which are neither convex nor concave w.r.t.
+$\mathbf{p}_{n}$.
+To address this problem, we employ the MM framework, where surrogate functions that upper bound $\widetilde{f}_{1}$ are constructed.
+Specifically, the construction of the surrogate functions is given by the following lemma.
+Lemma 1. [28] Considering a function $\widetilde{f}$ with $\mathbf{z}$ being the variable, its second-order Taylor expansion at $\mathbf{z}_{0}$ is given by
+$\widetilde{f}(\mathbf{z})\approx\widetilde{f}(\mathbf{z}_{0})$ $+(\nabla\widetilde{f}(\mathbf{z}_{0}))^{\mathrm{T}}(\mathbf{z}-\mathbf{z}_{0})$
+$+\frac{1}{2}(\mathbf{z}-\mathbf{z}_{0})^{\mathrm{T}}\nabla^{2}\widetilde{f}(\mathbf{z}_{0})(\mathbf{z}-\mathbf{z}_{0}),$ (34)
+where $\nabla\widetilde{f}(\mathbf{z})$ is the gradient vector of $\widetilde{f}(\mathbf{z})$, while $\nabla^{2}\widetilde{f}(\mathbf{z})$ represents the Hessian matrix of $\widetilde{f}(\mathbf{z})$.
+By selecting a parameter $\delta$ satisfying $\delta\mathbf{I}\succeq\nabla^{2}\widetilde{f}(\mathbf{z}_{0})$, the function $\widetilde{f}(\mathbf{z})$ can be upper-bounded by
+$\!\widetilde{f}(\mathbf{z})\leq\widetilde{f}(\mathbf{z}_{0})\!+\!(\nabla\widetilde{f}(\mathbf{z}_{0}))^{\mathrm{T}}(\mathbf{z}-\mathbf{z}_{0})\!+\!\frac{\delta}{2}\left\|(\mathbf{z}-\mathbf{z}_{0})\right\|^{2}.$ (35)
+In particular, $\delta$ is chosen as the Frobenius norm of $\nabla^{2}\widetilde{f}(\mathbf{z}_{0})$.
+According to Lemma 1, we separate $\mathbf{p}_{n}$ from the cosine functions and utilize the second-order Taylor expansion as the surrogate function.
+To construct the surrogate function, the gradient vector and the Hessian matrix of $\widetilde{f}_{1}(\mathbf{p}_{n})$ should be determined, which are given by the following theorem.
+Theorem 1. According to the linear property of derivative operation, the gradient of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the derivatives of its three components, which is expressed as
 $\nabla\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}}\end{bmatrix}^{\mathrm{T}}.$ (36)
-For the Hessian matrix of $\widetilde{f}_{1}$, it can also be given as
-$\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}\partial p_{y,n}}\\ \sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{y,n}}\end{bmatrix}.$ (40)
+For the Hessian matrix of $\widetilde{f}_{1}$, it is given by
+$\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})=\begin{bmatrix}\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{x,n}\partial p_{y,n}}\\ \sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p_{y,n}\partial p_{x,n}}&\sum_{i=1}^{3}\frac{\partial^{2}\widetilde{f}_{1,i}(\mathbf{p}_{n})}{\partial p^{2}_{y,n}}\end{bmatrix}$.
+(37)
 Specifically, there are 6 first-order derivative terms in (36) and 12 second-order derivative terms in (37), whose detailed expressions are left in the Appendix C for brevity.
 Proof: Please refer to Appendix C.
-According to Lemma 1 and Theorem 1, the equivalent objective function ultimately be given as
+According to Lemma 1 and Theorem 1, the relaxed objective function is given by
 $(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\!+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2},$ (38)
-where the constants are neglected and $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$.
-The $\mathbf{p}_{n}^{(t-1)}$ represents the output position of $n$-th element from the last round optimization.
-Apparently, the function (41) is a convex quadratic function with respect to $\mathbf{p}_{n}$.
-Apart from the objective function, the constraint (34c) remains non-convex in the current form.
-To deal with it, an effective approach is to relax the constraint (34c) with the first-order Taylor expansion.
-Specifically, the first-order Taylor expansion of the left term in (34c) can be given as
-$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}$
-$~{}~{}~{}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (42a)
+where the constants are neglected, while $\delta_{n}=\|\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n})\|^{2}_{\mathrm{F}}$.
+The $\mathbf{p}_{n}^{(t-1)}$ represents the position of the $n$-th element from the previous optimization result.
+Apparently, (38) is a convex quadratic function w.r.t.
+$\mathbf{p}_{n}$.
+Remark 4: The construction of (38) guarantees that the surrogate and the original function match in both value and gradient at the current point $\mathbf{p}_{n}^{(t-1)}$, ensuring local approximation.
+In particular, the approximation error can be characterized as
+error $\approx\frac{1}{2}\Delta\mathbf{p}_{n}^{\mathrm{T}}\left(\delta_{n}\mathbf{I}-\nabla^{2}\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)})\right)\Delta\mathbf{p}_{n},$ (39)
+where $\Delta\mathbf{p}_{n}=\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)}$.
+The upper-bounding strategy ensures that the surrogate becomes increasingly tight during convergence.
+Nevertheless, since the original problem is non-convex, the MM algorithm is only guaranteed to converge to a stationary point, which is generally a local optimum.
+Apart from the objective function, (14f) remains non-convex.
+To deal with it, we relax (14f) with the first-order Taylor expansion.
+Specifically, the first-order Taylor expansion of the left term in (14f) is given by
+$\|\mathbf{p}_{n}-\mathbf{p}_{n^{\prime}}\|_{2}$ $\geq\|\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}+\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (40a)
 $=\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}$ (40b)
-As a result, the constraint (34c) can be relaxed as
+As a result, (14f) is relaxed as
 $\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n,n^{\prime}\!=\!1:N,n\!\neq\!n^{\prime},$ (41)
-which is affine regarding $\mathbf{p}_{n}$. Therefore, the problem for optimizing the position of the $n$-th element is given as
-$\min_{\mathbf{p}_{n}}$ $~{}(\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)}))^{\mathrm{T}}(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}-\mathbf{p}_{n}^{(t-1)})\right\|^{2}$ (44a)
-$\mathrm{s.t.}$ $~{}\mathbf{p}_{n}\in\mathcal{A},$ (44b)
-$\!\frac{(\mathbf{p}_{n}^{(t\!-\!1)}\!-\!\mathbf{p}_{n^{\prime}})^{\mathrm{T}}\!(\mathbf{p}_{n}\!-\!\mathbf{p}_{n^{\prime}})}{\|\mathbf{p}_{n}^{(t-1)}\!-\!\mathbf{p}_{n^{\prime}}\|_{2}}\!\geq\!\Delta D,\forall n\!\neq\!n^{\prime}.$ (44c)
-The problem (44) is a typical convex quadratic problem, which can be solved by using quadprog [44, 45].
-Besides, we adopt circle packing scheme [46] for the initialization of fRIS element positions, where the initial positions are sufficiently separated.
+which is affine regarding $\mathbf{p}_{n}$.
+Notably, this method relaxes the non-convex feasible region to an inner half-space feasible region.
+Since part of the feasible region is discarded, the optimization may only find a suboptimal solution.
+Nonetheless, such approximation ensures algorithmic stability and is commonly accepted in MM-based optimizations.
+Therefore, the problem for optimizing the position of the $n$-th element is given by
+$\!\!\!\!\min_{\mathbf{p}_{n}}~\nabla\widetilde{f}_{1}(\mathbf{p}_{n}^{(t-1)})^{\mathrm{T}}(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})+\!\frac{\delta_{n}}{2}\left\|(\mathbf{p}_{n}\!-\!\mathbf{p}_{n}^{(t\!-\!1)})\right\|^{2}$ (42a)
+[…]
+which is a typical convex quadratic problem, and solved by using quadprog [37].
+Besides, we adopt the circle packing scheme [38] for the initialization of fRIS element positions, where the initial positions are sufficiently separated.
+This scheme guarantees minimum spacing and provides sufficient mobility space for subsequent iterations to avoid collision.
 The details of the proposed algorithm are summarized in Algorithm 1.
+Remark 5: Element repositioning may introduce additional overhead as compared to the traditional RIS.
+On the one hand, element displacement consumes energy, which mainly stems from the step motor drive or the port switching.
+On the other hand, element mobility adds computational complexity to position optimizations.
+Nonetheless, the extra spatial DoFs of fRIS allow the use of fewer elements, keeping the total computational burden reasonable.
 Algorithm 1 Optimization of fRIS element positions
-0:    Channel parameters, transmit signal, fRIS phase shift matrix, system parameters, desired reflective signal and communication symbols;
-1:  Initialize the element positions $\mathbf{p}^{(0)}$ with the circle packing scheme;
+0:   Channel parameters, transmit signal, fRIS phase shift matrix, reference signal, and communication symbols;
+1: Initialize $\mathbf{p}^{(0)}$ with the circle packing scheme;
 2:for $n=1\longrightarrow N$ do
@@
 III-F Convergence and Computational Complexity Analysis
-The proposed algorithm is summarized in Algorithm 2, which incorporates the solutions derived above.
-The algorithm iteratively optimizes the variables, where the process continues until the objective function value in (19) becomes stable.
+The proposed algorithm is summarized in Algorithm 2, which will terminate when the change of $\varepsilon$ falls below a threshold $\epsilon_{0}$ or the maximum iteration number $I_{\mathrm{max}}$ is reached.
 III-F1 Convergence Analysis
-In this part, we analyze that the convergence of the proposed Algorithm 2.
-First, the subproblem (20) is a unconstrained convex quadratic problem with a close-form solution, and thus $\omega^{t}$ update does not increase objective function value:
-$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (45)
-Updating $\mathbf{\Theta}^{t}$ in (23), no matter utilizing the SDR [39] or the manifold method [40], it was proved that the objective function value of (23) is non-increasing over iterations, such that
-$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (46)
-In the optimization of $\mathbf{W}$, quasi-Newton method guarantees that the update of $\overline{\mathbf{w}}$ is non-increasing [42], while equality constraint is satisfied if the penalty parameter becomes sufficiently large.
-As a result, the solution to (33) can satisfy the equality constraint, while the process (33) is non-increasing:
-$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t}\!,\mathbf{p}^{t}).$ (47)
-Proved in [25], the subproblem (44) is a non-decreasing process given the nature of its convexity and the inequalities of Taylor expansions, while the constrained optimization process can converge in the MM framework when the function is smooth objective function [47]:
-$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t+1})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t}).$ (48)
-Besides, the finite resources indicate that the objective function value is bounded and greater than zero. As a result, the Algorithm 2 can guarantee to converge.
+The convergence behavior of the proposed Algorithm 2 is analyzed as below.
+(i) Subproblem of symbol estimator optimization: The update of $\omega$ in (19) solves a convex quadratic problem, which guarantees $\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})$.
+(ii) Subproblem of phase shift optimization: The subproblem (22) involves the modulus-one constraint.
+It is solved via either SDR [42] or manifold optimization [49], where the SDR is emphatically shown as a non-increasing process and manifold optimization updates variables along the Riemannian gradient direction on the complex circle manifold.
+Both approaches ensure $\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{x}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})$.
+(iii) Subproblem of transmit waveform optimization: In the optimization of $\mathbf{x}$, the quasi-Newton method guarantees that the update of $\overline{\mathbf{x}}$ is non-increasing [3], while the equality constraint is satisfied if the penalty parameter becomes sufficiently large.
+As a result, the solution to (30) can satisfy the equality constraint, while the process is non-increasing: $\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{x}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{x}^{t}\!,\mathbf{p}^{t})$.
+(iv) Subproblem of element position optimization: The update of $\mathbf{p}$ is tackled via the MM framework, where a convex quadratic surrogate function is constructed.
+According to the convergence theory of MM methods [34], the surrogate ensures that the original objective is non-increasing, i.e., $\varepsilon(\mathbf{p}_{n}^{(i-1)})=g(\mathbf{p}_{n}^{(i-1)})+C\geq g(\mathbf{p}_{n}^{(i)})+C=\varepsilon(\mathbf{p}_{n}^{(i)})$, where $g(\cdot)$ is the surrogate function while $C$ is a constant.
+Besides, the inter-element spacing constraint is relaxed to a half space, which does not affect the monotonicity of the descent.
+Moreover, due to the non-negativity of MSEs and the finite transmit power, the overall objective function is lower-bounded by 0.
+Hence, the proposed Algorithm 2 generates a non-increasing and bounded sequence of objective values, and converges to a stationary point that is at least a local optimum.
 III-F2 Computational Complexity Analysis
 In this part, the computational complexity is analyzed.
-  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving the problem (16) with SDR and the eigenvalue decomposition.
+  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving (15) with SDR and the eigenvalue decomposition.
 Given a solution accuracy $\epsilon_{1}$, the computational complexity of SDR is $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$ [21], while the eigenvalue decomposition has the computational complexity of $\mathcal{O}\left(N^{3}\right)$.
-  • The optimal symbol estimator can be calculated via a close-form expression, whose computational complexity stems from matrix multiplications.
+  • The optimal symbol estimator can be calculated via a closed-form expression, whose computational complexity stems from matrix multiplications.
 Obtaining the numerator in (20) has the computation complexity of $\mathcal{O}\left(KN+N^{2}+MN+M\right)$, while calculating the denominator has the computation complexity of $\mathcal{O}\left(2(MK+NK+N^{2})+M\right)$.
-  • When optimize the fRIS phase shift, the computational complexity of calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$.
+  • When optimizing the fRIS phase shift, the computational complexity of the calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$.
 Besides, the SDR approach has the computational complexity of $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{2})\right)$ with $\epsilon_{2}$ being the solution accuracy, while that of the manifold optimization is $\mathcal{O}\left(N^{1.5}\right)$ [49].
-  • In the optimization of transmit beamformer design, the computational complexity of matrix multiplication is $\mathcal{O}\left(M^{2}K^{2}\!+\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!+\!K^{2}\right)$, while the design of the transmit beamformer with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$ with $\epsilon_{3}$ being the convergence accuracy [48, 49].
-  • When optimize the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$.
+  • In the optimization of transmit waveform design, the computational complexity of matrix multiplication is $\mathcal{O}\left(\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!\right)$, while the design of the transmit waveform with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$, where $\epsilon_{3}$ is the convergence accuracy [14, 11].
+  • When optimizing the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$.
 Additionally, the computational complexity of solving the problem (42) is $\mathcal{O}\left(\left(N^{1.5}\log(1/\epsilon_{4})\right)\right)$ [23] with accuracy $\epsilon_{4}$ for the interior-point method.
-As a result, the total computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$, neglecting the lower-order terms.
+In summary, the computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon)\right)$, neglecting the lower-order terms.
 Algorithm 2 Proposed alternating optimization algorithm
-0:    Channel parameters, noise power, transmit power budget, ideal sensing beampattern, weighting coefficient, users number, fRIS planar region size, and the minimum inter-element spacing requirement.
-1:  Initialize all the variables;
-2:  while no convergence of $\varepsilon$ do
-3:     Calculate correlation matrix for sensing-only by (16) and obtain desired reflective signal $\mathbf{s}_{r}$ via (17);
+0:   Channel parameters, noise power, transmit power, ideal sensing beampattern, weighting coefficient, fRIS size, and the minimum inter-element spacing.
+1: Initialize all the variables, and $i=0$;
+2:while ${|\varepsilon^{(i+1)}-\varepsilon^{(i)}|}/{\varepsilon^{(i)}}>\epsilon_{0}$ or the maximum iteration number is not reached, do
+3:  Calculate correlation matrix by (15) and obtain reference signal $\mathbf{s}_{r}$ via (16);
 4:  Update communication symbol estimator by (20);
 5:  Optimize $\boldsymbol{\Theta}$ by solving problem (22);
-6:     Update ${\mathbf{W}}$ by implementing the ALM process in (33);
+6:  Update $\mathbf{x}$} by implementing the ALM process in (30);
 7:  Optimize fRIS element positions via Algorithm 1;
-8:  end while
-9:  return  Optimized variables $\mathbf{W}$, $\boldsymbol{\Theta}$, $\mathbf{p}$, $\omega$, $\beta$ and the objective function value.
+8:  Update iteration index $i=i+1$;
+9:end while
+10:return Optimized variables and joint MSE.
 IV Simulation Results
-This section provides numerical results to validate the effectiveness of the fRIS-aided ISAC system and the proposed algorithm.
-Unless stated otherwise, the simulation parameters are set as follows.
-The center frequency of this system is $2.4$GHz, corresponding to a wavelength of $\lambda\!=\!0.125$m.
-The ISAC BS is located at ($3$m, 00m, 00m), and equipped with $M\!=\!8$ antennas.
-The transmitter has a maximum transmit power of $10$dBm.
-The fRIS, consisting of $16$ elements ($N_{x}\!=\!N_{y}=4$), is positioned at (00m, $3$m, $3$m), with a size of $A\!\times\!A$ where $A\!=\!4\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$.
-Additionally, we consider $K\!\!=\!\!4$ communication users and $T\!\!=\!\!3$ targets.
-The users are randomly distributed within an area centered at ($30$m, $100$m, 00m) with a radius of $10$m.
-As observed from fRIS, the targets’ (azimuth, elevate) angles are ($-60^{\circ}$, $0^{\circ}$), ($10^{\circ}$, $0^{\circ}$) and ($55^{\circ}$, $0^{\circ}$), respectively.
-The noise power of the users are set as $\sigma_{0}^{2}\!=\!-60$dBm and QPSK is used for modulation.
-The weighting coefficient $\alpha$ for balancing the communication and sensing performance is $0.5$.
-Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-10$dB is the path loss exponent and dist denotes the transmitter-receiver distance [39].
+This section provides numerical results to validate the effectiveness of the system and the proposed algorithm.
+We set the default parameters as follows.
+The center frequency of this system is 2.42.4GHz.
+The ISAC BS is located at ($3$, 0, 0)m, and equipped with M=8M\!=\!8 antennas.
+The transmitter has maximum transmit power of 1010dBm.
+The fRIS, consisting of $25$ elements ($N_{x}\!=\!N_{y}=5$), is positioned at (0, 33, 33)m, with a size of $A\!\times\!A$ where $A\!=\!5\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$.
+Additionally, we consider K=4K\!\!=\!\!4 users and T=3T\!\!=\!\!3 targets.
+The users are randomly distributed within an area centered at ($50$, $100$, 0)m with a radius of 1010m.
+Observed from fRIS, the targets’ (azimuth, elevation) angles are ($20^{\circ}$, $0^{\circ}$), ($5^{\circ}$, $0^{\circ}$), and ($30^{\circ}$, $0^{\circ}$), respectively.
+The noise power is set as $\sigma_{0}^{2}\!=\!-60$dBm, and QPSK is used for modulation.
+The weighting coefficient $\alpha$ is $0.5$.
+Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-3$dB is the path loss, and dist denotes the transmitter-receiver distance [42].
 The penalty factor for ALM and the convergence threshold are set as $1$ and $10^{-5}$, respectively.
-All simulations are averaged over $100$ Monte Carlo trials.
-Figure 2: Convergence behaviors of the proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step.
-Figure 3: The effect of RIS element number: the objective function value $\varepsilon$ versus the element number.
+All simulation results are averaged over $100$ Monte Carlo trials.
+In addition, $\mathbf{x}$, $\boldsymbol{\Theta}$, and $\omega$ are randomly initialized under their respective constraints, while the fRIS element positions $\mathbf{p}$ are initialized as stated in Section III-E.
+(a)                                     (b)
+Figure 2: Convergence behaviors: (a) The proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step.
+(b) Fluctuations with a fixed channel realization and random initialization parameters.
 The performance of the proposed design is evaluated against the following benchmark schemes:
-  • Conventional RIS-aided ISAC system (Conven): The positions of RIS elements are fixed and uniformly distributed across the fRIS plane.
+  • Conven (SameNum): This scheme is an ISAC aided by fixed-element RIS, where the RIS has the same element number as fRIS (sparse).
 The rest of the beamforming designs follow the same procedures as proposed.
-  • Discrete position selection (DPS): The moving region of the elements is quantized into discrete grids of equal spacing, where greedy algorithm is employed for the position selection.
-Specifically, the position of each element is selected sequentially in the discrete position set, while the positions of the other elements are kept unchanged.
-  • Random RIS element deployment (Rand): The positions of RIS elements remain fixed for each channel realization but randomly vary across realizations.
-The optimization of other components is performed as the proposed scheme.
-Figure 4: Beampattern at the fRIS: beam gain (dB) versus spatial angle (degree).
-(a) The beampattern when bias to the communication ($\alpha=0.1$).
+  • Conven (HWL): The ISAC system is aided by a fixed-element RIS, where the elements have dense half-wavelength spacing in the same surface size.
+The rest of the optimization procedures are identical to the proposed.
+  • DPS: In this scheme, the fRIS element position is discrete, where the fRIS is quantized into discrete grids of equal spacing.
+Specifically, the continuous position is projected to the nearest position.
+Figure 3: Beampattern at the fRIS: beam gain (dB) versus spatial angle ().
+(a) The beampattern when biased to the communication ($\alpha=0.1$).
 (b) The beampattern when no bias ($\alpha=0.5$).
-(c) The beampattern when bias to the sensing ($\alpha=0.9$).
+(c) The beampattern when biased to the sensing ($\alpha=0.9$).
 (d) The beampattern of different approaches when $\alpha=0.5$.
-TABLE I: Comparisons in terms of Average Performance Value
-and Time Complexity per round
-Scheme Objective Function Value Elapsed Time (s)
-Proposed 0.378 313.52
-($N=25$)
-Conven 0.394 934.01
-($N=64$)
-Conven 0.432 372.46
-($N=49$)
-In Fig.2, we examine the convergence performance of our proposed algorithm for the fRIS-assisted ISAC system and compare it with the benchmark scenarios.
-As expected, the proposed approach demonstrates convergence, with the objective function value decreasing rapidly and stabilizing at $0.57$ after $15$ iterations.
-Meanwhile, it can be observed that the joint MSE of the proposed fRIS-aided ISAC case outperforms the baselines, especially compared to the conventional RIS-aided ISAC case.
-This improvement can be attributed to the continuous and flexible element position optimization, which provides additional DoFs in the spatial domain.
-Fig.3 illustrates the relationship between the number of fRIS elements ($N$) and the objective function value ($\varepsilon$), with $N$ varying from $16$ to $64$.
-As observed, the objective function value of the proposed scheme and the baselines all decrease as $N$ increases, while the proposed scheme achieves the best performance among all.
-In addition, in Table I, we compare the average objective function value $\varepsilon$ and the time complexity of the proposed scheme with those of the conventional RIS-aided ISAC system.
-Specifically, we select the proposed scheme with $N=25$, the “Conven” scheme with $N=49$ and $N=64$, where either the objective function value or the elapsed time has similar performance.
-When compared with the “Conven” scheme with $N=64$, the proposed scheme with $N=25$ can achieve slightly better objective function value yet requiring $1/3$ of the elapsed time.
-When compared with the “Conven” scheme with $N=49$, the proposed scheme with $N=25$ exhibits comparable elapsed time but significantly better objective function value performance.
-These results demonstrate that employing the fRIS not only compensates for performance losses caused by reducing RIS elements but also alleviates the complexity of high-dimensional optimization.
-To evaluate the system’s sensing performance, we further investigate the fRIS beampattern under different weighting coefficients, and compare with the ideal case and the baselines in Fig.4.
-It is worth noting that the targets are assumed to lie at an angle of $0^{\circ}$ in elevate, so that the beampatterns of the cross section at $0^{\circ}$ in elevate are plotted for clearer observations and easier comparisons.
-In Fig.4 - Fig.4, the practical, sensing-only and ideal beampatterns are illustrated for the weighted coefficients $\alpha=0.1,0.5\text{ and }0.9$, respectively.
-It can be observed that beams are consistently formed in the target directions across all $\alpha$ values.
-However, in the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels.
-In contrast, when $\alpha=0.5$ and $\alpha=0.9$ where the system is increasingly biased towards the sensing, the practical beams gradually align with the sensing-only case.
-On the other hand, Fig.4 compares the beampattern obtained from different approaches when $\alpha=0.5$.
-It can be found that all the beampatterns have formed peaks in the target directions, but the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the ”DPS” and ”Rand” cases.
-Figure 5: Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients.
-Figure 6: Average BER (dB) versus different receive noise power (dBm) with different approaches.
-In addition, we evaluate the communication performance of the system under different settings.
-Fig.5 depicts the average bit error rate (BER) of the proposed scheme versus various noise power levels, ranging from $-40$dBm to $-80$dBm.
-As expected, the average BER decreases when the noise power at the receivers decreases, and QPSK modulation exhibits lower BER when compared to 16QAM.
-Moreover, when $\alpha$ decreases, the BER further improves as the system becomes more communication-biased.
-Notably, at a noise power of $-80$dBm and a weighting coefficient of $\alpha=0.9$, the BER can drop to approximately $-28$dB with QPSK modulation.
+Figure 4: Beampatterns (dB) of different system configurations: (a) Angles Separation: $5^{\circ}$, Element number: $16$, Size: $8\lambda$.
+(b) Angles Separation: $10^{\circ}$, Element number: $16$, Size: $8\lambda$.
+(c) Angles Separation: $5^{\circ}$, Element number: $16$, Size: $12\lambda$.
+(d) Angles Separation: $5^{\circ}$, Element number: $64$, Size: $8\lambda$.
+(e) Angles Separation: $10^{\circ}$, Element number: $64$, Size: $8\lambda$.
+(f) Angles Separation: $5^{\circ}$, Element number: $64$, Size: $12\lambda$.
+TABLE I: Confirmation of the joint effect of position and phase optimization of fRIS
+Objective function Commun. Sensing
+Value MSE MSE
+Joint 0.27 0.23 0.31
+optimization
+Position 0.79 1.1 0.48
+only
+Phase shift 0.57 0.67 0.47
+only
+No phase shift or 2.07 3.51 0.63
+position optimization
+In Fig.
+2(a), we examine the convergence performance of the proposed algorithm for the fRIS-aided ISAC system and compare it with the benchmarks.
+As expected, the proposed approach demonstrates rapid convergence, stabilizing after $20$ iterations.
+Besides, the MSE of the proposed scheme outperforms the baselines under the same element number, while yielding comparable performance as the “Conven (HWL)” case.
+This is because the continuous and flexible element position optimization provides additional spatial DoFs.
+In addition, the DoFs of element position optimization are validated in Table I.
+As compared to position-only/phase-only optimization, the joint optimization outperforms the individual optimization schemes, implying that repositioning does not overlap with the existing optimization dimension.
+To assess the impact of initialization, we conduct $100$ trials under a fixed channel realization and different initialization.
+As shown in Fig.
+2(b), the convergence exhibits small variation, with a mean of $0.1461$, a variance of $0.0065$, and a range of [$0.0577$, $0.4922$].
+This indicates good robustness and stable convergence behavior.
+TABLE II: Comparisons in terms of Average Performance Value
+and Time Complexity per iteration
+Scheme Objective Function Value Elapsed Time
+(s/iteration)
+Proposed: 0.27 24.6
+$N=25$
+Conven (SameNum): 0.57 1.33
+$N=25$
+Conven (HWL): 0.32 47.75
+$N=81$
+To further validate the advantages of fRIS, we compare the average objective function value and the time complexity of the proposed scheme with those of the conventional RIS-aided schemes in Table II.
+Specifically, in the “Conven (SameNum)” scheme, the conventional RIS has the same element number as fRIS ($N=25$), while in the “Conven (HWL)” scheme, the elements have fixed half-wavelength spacing under the same aperture, where the element number is $N=81$.
+It is seen that the proposed scheme achieves better results compared to “Conven (SameNum)” due to additional DoFs.
+When compared with the “Conven (HWL)” scheme, the proposed scheme exhibits comparable performance but significantly less elapsed time per iteration.
+These results demonstrate that employing the fRIS can compensate for performance losses caused by reducing RIS elements, and alleviate the complexity of high-dimensional optimization.
+For the sensing performance evaluation, we demonstrate the fRIS beampattern under different weighting coefficients ($\alpha=0.1,0.5,0.9$), and compare with the ideal case and the reference beampattern in Figs.
+3-3.
+It is observed that beams are consistently formed in the target directions.
+In the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels.
+In contrast, when $\alpha=0.5,0.9$ where the system is increasingly biased towards the sensing, the formed beams gradually align with the reference beampattern.
+While a discrepancy in beampattern energy exists between the “Proposed” and “Sensing-only” cases, it arises from the joint optimization process, where the inherent tradeoff introduces a mismatch between the covariance of the actual signal $\mathbf{v}$ and that of the reference signal $\mathbf{s}_{r}$.
+Consequently, the realized signal may differ from $\mathbf{s}_{r}$ in both shape and power.
+In addition, since $\mathbf{s}_{r}$ is a virtual design target and not physically transmitted, this mismatch does not violate energy conservation.
 On the other hand, Fig.
-6 compares the average BER of the proposed method to the benchmarks.
+3 compares the beampattern of different approaches.
+It is observed that all the beampatterns have formed peaks in the target directions.
+Meanwhile, the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the “DPS” cases.
+To further evaluate the sidelobe levels and the energy leakage, we demonstrate integrated sidelobe to mainlobe ratio (ISMR) [4] in Table III.
+Specifically, ISMR is computed as the ratio of the integrated power in the sidelobe region to that in the mainlobe region.
+Given the target directions $\left[-20^{\circ},5^{\circ},30^{\circ}\right]$, we set the mainlobe regions as $\left[-30^{\circ},-10^{\circ}\right]\cup\left[-5^{\circ},15^{\circ}\right]\cup\left[20^{\circ},40^{\circ}\right]$.
+It is observed that higher $\alpha$ yields lower ISMR, demonstrating better sidelobe suppression, while the proposed fRIS-aided design consistently achieves lower ISMR than the conventional RIS scheme with the same amount of element.
+For further sidelobe suppression, we may consider incorporating ISMR into optimization.
+Such enhancement is non-trivial and worth exploring in future work.
+Under closely spaced targets, we simulate the beampatterns when two targets are placed at $0^{\circ}$ and another at $5^{\circ}$ or $10^{\circ}$.
+The aperture size of fRIS is set as $A\!\in\!\{8\lambda,12\lambda\}$ and the number of elements is set as $N\!\in\!\{16,64\}$.
+Fig.
+4 reveals that larger angular separation improves beam distinguishability, while increasing aperture $A$ narrows the mainlobes.
+For instance, while $A\!=\!8\lambda$ leads to merged beams at $5^{\circ}$ separation, $A\!=\!12\lambda$ achieves separated peaks.
+Additionally, increasing $N$ can suppress the sidelobe and enhance the mainlobe gain.
+TABLE III: ISMR under different settings.
+$\alpha=0.1$ $\alpha=0.5$ (Default) $\alpha=0.9$
+ISMR (dB) 1.49 -0.22 -0.51
+Conven (SameNum): Conven (HWL): DPS
+$N=25$ $N=81$
+ISMR (dB) 2.65 -0.58 -0.18
+(a)                                     (b)
+Figure 5: (a) Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients.
+(b) Average BER (dB) versus different receive noise power (dBm) with different approaches.
+In addition, we evaluate the communication bit error rate (BER) of the system under different settings.
+Fig.
+5(a) depicts the BER of the proposed scheme versus various noise power levels.
+As expected, the average BER decreases when the noise power decreases, and QPSK modulation exhibits lower BER when compared to higher-order modulations.
+Moreover, when $\alpha$ decreases, the BER improves as the system becomes more communication-biased.
+On the other hand, Fig.
+5(b) compares the average BER of the proposed method to the benchmarks.
 As shown, all cases exhibit a decreasing BER trend as noise power level reduces, while the proposed method consistently achieves a lower BER compared to the baselines.
-Figure 7: The effect of the fRIS region size: the objective value versus the normalized region size $A/\lambda$.
-In Fig.7, the impact of the normalized region size $A/\lambda$ on the proposed scheme is illustrated, where $A$ varies from $2\lambda$ to $6\lambda$.
-The results show that the objective function value $\varepsilon$ decreases as the region size increases, which is expected since the larger size of the fRIS provides more DoFs for the mobility of the elements.
-On the other hand, the benchmarks of “DPS” and “Rand” cases demonstrate a decreasing trend with increasing region size, while the “Conven” case has little change as the region size varies.
-Compared with the benchmarks, the proposed scheme and the optimization algorithm achieve superior performance across all region sizes, which is consistent to the previous simulation results.
+In Fig.
+6, the impact of the fRIS size on BER is evaluated, where we find that the larger fRIS aperture size yields lower BER.
+This is because increasing $A$ can form narrower beams, which increase the received energy for users.
+Figure 6: Average BER (dB) versus different receive noise power (dBm) under different fRIS sizes.
 V Conclusions
-This paper proposed a fRIS-aided ISAC system to achieve multi-user communications and multi-target sensing by exploiting the beamforming and the fRIS element position optimization.
-We formulated a joint communication symbol estimation MSE and sensing beampattern mismatch minimization problem subject to transmit power budget, constant-modulus constraint and the element position requirements of the fRIS.
-An AM-based algorithm was developed, where each subproblem utilized the techniques of derivatives, SDR, ALM and MM optimization.
-Simulation results validated that the fRIS can significantly reduce the sensing beampattern mismatch and communication symbol estimation error in the ISAC systems, as compared to the conventional RIS.
+This paper proposed a fRIS-aided ISAC system to achieve multi-user communication and multi-target sensing, where a joint communication MSE and sensing beampattern mismatch minimization problem was formulated.
+To address the problem, an AM-based algorithm was developed, which involved various techniques.
+Simulation results validated that the fRIS significantly reduces the sensing beampattern mismatch and communication BER.
+Future research may explore new fRIS architectures in ISAC systems, such as active or beyond diagonal fRIS designs.
+In addition, it is interesting to study the modeling of fRIS under rich-scattering environments.
 Appendix A Derivation of the function $\widetilde{f}_{0}$ in (III-E)
-To simplify the objective function (34a), we first expand the two square norms, and it is given as
+To simplify (31a), we first expand the two square norms, which is given by
 $\!\!\!\alpha\cdot\left(\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}-2\mathfrak{R}\left\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right)\!+\!(1\!-\!\alpha)\cdot$
 $\!\!\!\!\left(\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\!\!-\!\!2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}\right).$ (43)
-It is noteworthy that $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
-Observing the first term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is actually unrelated to the variable $\mathbf{p}$ due to the following reason.
-Given the property of $\boldsymbol{\Theta}$ that $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (10), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$.
-Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (8), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of element positions.
-As a result, this term eventually becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
-For the second term $-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can also be substituted, yielding $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$.
-Leveraging the cyclic property of trace, it can be re-expressed as $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
-Focusing on the third term, it can be rewritten as
+Here, $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
+Observing the term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is unrelated to $\mathbf{p}$ due to the following reason.
+Given $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (9), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$.
+Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (7), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of $\mathbf{p}$.
+Therefore, this term becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
+For the term $2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can be substituted, yielding $2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$.
+Leveraging the cyclic property of trace, it is re-expressed as $2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
+Focusing on the third term, it is rewritten as
 $\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{H}_{rc}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$ (44a)
@@
 $\overset{\mathrm{(c)}}{\longrightarrow}$ $c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p}).$ (44d)
-In operation (a), the expression of $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms for simplicity.
-In operation (b), it is derived based on the matrix structure in (14a) and (14b).
+In (a), the expression of $\mathbf{G}$ and $\mathbf{H}_{rc}$ are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms.
+In (b), it is derived based on the matrix structure in (13a) and (13b).
 The operation (c) leverages the diagonal property of $\boldsymbol{\Theta}$, where we denote $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$.
-As for the last term in (A), it can be transformed as follows:
+For the last term in (A), it is transformed as follows:
 $-2\omega\mathfrak{R}\left\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\right\}$ (45a)
@@
 where (a) is the substitution operation, while (b) utilizes the cyclic property of trace and the diagonal property of $\boldsymbol{\Theta}$.
-We also denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
+In addition, we denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
 By integrating all the simplified terms, the objective function can be expressed as (III-E). This ends the derivation.
 Appendix B Derivation of the function $\widetilde{f}_{1}$ in (III-E)
-For the function (III-E), it can be further simplified as follows if one considers the position of $n$-th element position $\mathbf{p}_{n}$ while fixing others.
-First, examining the first term $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which can be expressed as
+For (III-E), it can be simplified if we consider the position optimization of the $n$-th element $\mathbf{p}_{n}$ while fixing others.
+First, examining the first term $2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which is expressed as
 $\sum_{n=1}^{N}-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (46a)
 $\overset{\mathrm{(a)}}{\longrightarrow}$ $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\left\{[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}e^{\jmath\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})}\right\}$ (46b)
-[…]
-In particular, $[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$.
+${\longrightarrow}$ $-\nu_{n}\cos\left(\xi_{n}+\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})\right).$ (46c)
+In particular, $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is the $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$.
 Additionally, $\nu_{n}=2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is a variable-independent value.
-The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, which is given as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$.
-Moreover, in the operation (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
-For the second term, one can rewrite it in element-wise manner, i.e.,
+The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, i.e., $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$.
+Moreover, in (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
+For the second term, it can be rewritten in an element-wise manner, i.e.,
 $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\widetilde{\mathbf{A}}_{c,k}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})$ (47a)
@@
 where $d_{c,k,n}(\mathbf{p}_{n})=p_{x,n}\sin(\phi_{c,k})\cos(\psi_{c,k})+p_{y,n}\sin(\psi_{c,k})$.
-Apparently, when $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$.
-On the other hand, it can be found that the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering $n$-th element (i.e.
-$i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (53c) can be reduced to
+When $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$.
+Besides, the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering the $n$-th element (i.e.
+$i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (47c) is reduced to
 $(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\zeta_{k}\sum_{i=1,i\neq n}^{N}2\mathfrak{R}\left\{\left[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}\right]_{i,n}\right$.\!\!\!
 $\!\!\left.\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}\cdot e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}.$ (48)
-Notably, when fixing the positions of other elements, the term $[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$.
+Notably, when fixing the positions of other elements, the term $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$.
 We denote the phase of $\widetilde{c}_{i,n,k}$ as $\widetilde{\xi}_{i,n,k}=\angle\widetilde{c}_{i,n,k}$ while $\widetilde{c}_{i,n,k}$ has an amplitude of $1$.
-As a result, (B) can further be written as
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!2\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (55a)
-$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (55b)
-where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$ for simplicity.
+As a result, (B) can be written as
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!2(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!\!\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (49a)
+$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}\zeta_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (49b)
+where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}=2(1\!-\!\alpha)c_{0}$ for simplicity.
 Focusing on the last term, it can be transformed as follows:
 $-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{\mathrm{tr}\left(\mathbf{S}\mathbf{A}_{rc}^{\mathrm{H}}(\mathbf{p})\widetilde{\mathbf{A}}_{r}(\mathbf{p})\right)\right\}$ (50a)
-$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (56b)
-$\longrightarrow$ $\sum_{k=1}^{K}-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (56c)
-$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (56d)
+$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-\overline{\nu}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (50b)
+$\longrightarrow$ $\sum_{k=1}^{K}-\overline{\nu}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (50c)
+$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}~\overline{\rho}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (50d)
 where $s_{n,k}=[\mathbf{S}]_{n,k}$ with the amplitude of $\overline{\rho}_{n,k}=|s_{n,k}|$ and the phase of $\overline{\xi}_{n,k}=\angle s_{n,k}$.
-We also denote $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$ for brevity.
+We also denote $\overline{\nu}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}$ for brevity.
 This completes the derivation.
@@
 Given that $\widetilde{f}_{1}(\mathbf{p}_{n})$ is the sum of the sub-functions $\widetilde{f}_{1,1}$, $\widetilde{f}_{1,2}$, $\widetilde{f}_{1,3}$ shown in (III-E), the derivative of $\widetilde{f}_{1}(\mathbf{p}_{n})$ is equal to the respective derivatives of the sub-functions.
-For the first sub-function $\widetilde{f}_{1,1}$, the first-order derivative with respect to $p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear.
-As a result, the can be handled by the chain rule of derivative.
+For the sub-function $\widetilde{f}_{1,1}$, the first-order derivative w.r.t.
+$p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear.
+Therefore, we can adopt the chain rule of derivatives.
 First, we derive the the first-order derivatives of $\widetilde{f}_{1,1}$:
-$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,x}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57a)
-$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,y}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57b)
-where $\kappa_{r,x}=\sin(\phi_{r})\cos(\psi_{r})$ and $\kappa_{r,y}=\sin(\psi_{r})$. Based on (57), the second-order derivatives can be derived as
-$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58a)
-$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\!$ $=\!\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}$
-$=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}\kappa_{r,y}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58b)
-$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,y}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right).$ (58c)
-For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can also be derived as
-$\!\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59a)
-$\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59b)
-$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59c)
-$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}$
-$\!=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59d)
-$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59e)
-where $\Delta\widetilde{\kappa}_{x,k}=\kappa_{r,x}-\kappa_{c,x,k}$ and $\kappa_{c,x,k}=\sin(\phi_{c,k})\cos(\psi_{c,k})$, while $\Delta\widetilde{\kappa}_{y,k}=\kappa_{r,y}-\kappa_{c,y,k}$ and $\kappa_{c,y,k}=\sin(\psi_{c,k})$.
-Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given as
-$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60a)
-$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60b)
-$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60c)
-$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}$
-$=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60d)
-$\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right).$ (60e)
-By substituting the results of the first-order and second-order derivatives to (39) and (40), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained.
-This finishes the derivations.
+$\left[\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\right]^{\mathrm{T}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\boldsymbol{\kappa}_{r}\sin\!\left(\tau_{n}\right),$ (51a)
+where $\boldsymbol{\kappa}_{r}=[\kappa_{r,x},\kappa_{r,x}]^{\mathrm{T}}=[\sin(\phi_{r})\cos(\psi_{r}),\sin(\psi_{r})]^{\mathrm{T}}$ and $\tau_{n}=\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}$.
+Based on (51), the second-order derivatives are derived as
+$\begin{bmatrix}\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}&\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}&\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\end{bmatrix}=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\cos\!\left(\tau_{n}\right)\boldsymbol{\kappa}_{r}\boldsymbol{\kappa}_{r}^{\mathrm{T}}.$ (52)
+For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can be derived as
+$\!\!\!\left[\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\right]=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\boldsymbol{\widetilde{\kappa}}_{k}\sin\!\left(\widetilde{\tau}_{i,k,n}\right)$ (53a)
+$\!\!\!\begin{bmatrix}\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}&\!\!\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}&\!\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\end{bmatrix}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\mathbf{K}_{k}\cos\left(\widetilde{\tau}_{i,k,n}\right)\!,$ (53b)
+where $\widetilde{\tau}_{i,k,n}=\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}$, $\Delta\mathbf{K}_{k}=\Delta\boldsymbol{\widetilde{\kappa}}_{k}\Delta\boldsymbol{\widetilde{\kappa}}_{k}^{\mathrm{T}}$, $\Delta\boldsymbol{\widetilde{\kappa}}_{k}=\boldsymbol{\kappa}_{r}-\boldsymbol{\kappa}_{c,k}$ and $\boldsymbol{\kappa}_{c,k}=[\sin(\phi_{c,k})\cos(\psi_{c,k}),\sin(\psi_{c,k})]^{\mathrm{T}}$.
+Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given by
+$\!\!\!\!\!\left[\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\right]=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\boldsymbol{\widetilde{\kappa}}_{k}\sin\!\left(\overline{\tau}_{n,k}\right)$ (54a)
+$\!\!\!\!\!\begin{bmatrix}\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}^{2}}&\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}\partial p_{x,n}}&\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,3}}{\partial p_{y,n}^{2}}\end{bmatrix}\!\!=\!\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\mathbf{K}_{k}\cos\!\left(\overline{\tau}_{n,k}\right)\!,$ (54b)
+where $\overline{\tau}_{n,k}=\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}$.
+By substituting the above results to (36) and (37), the corresponding gradient and the Hessian matrix of $\widetilde{f}_{1}$ will be obtained. This finishes the derivations.